### v1

# Well-Posed Geometric Boundary data in General Relativity, III: Conformal-volume data

Abstract.
In this third work in a series, we prove the local-in-time well-posedness of the IBVP for the vacuum Einstein equations in general relativity with twisted DIrichlet boundary conditions on a finite timelike boundary. The boundary conditions consist of specification of the pointwise conformal class of the boundary metric, together with a scalar density involving a combination of the volume form of the bulk metric restricted to the boundary together with the volume form of the boundary metric itself.
1. Introduction
This work is the third in a series on the initial boundary value problem (IBVP) in general relativity. As in Parts I, II, [2, 3], the setting is the study of globally hyperbolic vacuum solutions of the Einstein equations
(1.1) $\mathrm{Ric}_{g}=0,$
on a finite domain $M\simeq I\times S$, where the Cauchy surface $S$ is a compact, $n$-dimensional, ($n\geq 2$), manifold with boundary $\partial S=\Sigma$ and ${\mathcal{C}}=I\times\Sigma$ is a timelike boundary ${\mathcal{C}}$ of $M$.
Consider the space $\mathbb{E}$ of all such vacuum solutions with a fixed topology $M$. Let $\mathrm{Diff}_{0}(M)$ be the group of diffeomorphisms $\varphi:M\rightarrow M$, isotopic to the identity and equal to the identity on $S\cup{\mathcal{C}}$. The group $\mathrm{Diff}_{0}(M)$ acts naturally on $\mathbb{E}$ and the quotient space ${\mathcal{E}}=\mathbb{E}/{\rm Diff}_{0}(M)$ is the moduli space of vacuum solutions, closely related to the covariant phase space of vacuum Einstein metrics. For simplicity, all data here and below are assumed to be $C^{\infty}$ smooth; cf. Remark 6.2 for extension to function spaces of finite differentiability.
The primary issue here is to find effective descriptions or parametrizations of ${\mathcal{E}}$ in terms of initial data on $S$ and boundary data on ${\mathcal{C}}$. Let then $\mathcal{I}_{0}$ denote the space of vacuum initial data, i.e. pairs $(\gamma_{S},\kappa)$ consisting of a Riemannian metric $\gamma_{S}$ on $S$ and a symmetric bilinear form $\kappa$ on $S$, satisfying the vacuum Einstein constraint equations, (discussed in detail later). In [2], the space of boundary data consisted of the space of conformal classes of globally hyperbolic Lorentz metrics on ${\mathcal{C}}$ and the space of scalar functions on ${\mathcal{C}}$ representing the mean curvature of the boundary ${\mathcal{C}}$ in $(M,g)$. In [3], we considered the space of Dirichlet boundary data, i.e.  globally hyperbolic Lorentz metrics on ${\mathcal{C}}$. In this work, we again choose the space of boundary data
(1.2) ${\mathcal{B}}={\rm Conf}({\mathcal{C}})\times C_{+}^{\infty}({\mathcal{C}}),$
to be the space of pointwise conformal classes of globally hyperbolic Lorentz metrics $\gamma_{{\mathcal{C}}}$ on ${\mathcal{C}}$ times the space of smooth positive scalar functions on ${\mathcal{C}}$. The boundary data of $(M,g)$ are now given by the pair
(1.3) $([g_{{\mathcal{C}}}],\mu_{g}),$
consisting of the conformal class $[g_{{\mathcal{C}}}]$ of the induced boundary metric, together with the scalar
(1.4) $\mu_{g}=dv_{g}(dv_{g_{{\mathcal{C}}}})^{-2/n}.$
Here $dv_{g}$ is the volume form of the ambient metric $g$ on $M$ restricted to ${\mathcal{C}}$ and $dv_{g_{{\mathcal{C}}}}$ is the volume form of the boundary metric $g_{{\mathcal{C}}}$ on ${\mathcal{C}}$. The product $\mu_{g}$ is viewed as a scalar density with respect to an arbitrarily chosen background volume form on $M$. The particular form of $\mu_{g}$ in (1.4) is chosen since it interacts well with the choice of gauge used here - the harmonic or wave coordinate gauge. Different choices of gauge may lead to other expressions for the scalar $\mu_{g}$, but this will not be pursued further here.
Note that the pair of boundary data $([g_{{\mathcal{C}}}],dv_{g_{{\mathcal{C}}}})$, when taken together, is exactly Dirichlet boundary data $g_{{\mathcal{C}}}$ studied in [3] - which was proved to be ill-posed (i.e. not well-posed) in general, i.e. without further conditions. Thus the boundary data (1.3) can be viewed as a modification of Dirichlet boundary data, involving a twisting by the volume form of $g$. The boundary data (1.3) involves no derivatives of the metric at ${\mathcal{C}}$.
For ${\mathcal{B}}$ as in (1.2), there is a natural map to initial and boundary data given by
(1.5) $\begin{split}&\Phi:\mathbb{E}\rightarrow\mathcal{I}_{0}\times_{c}{\mathcal{B}},\\ &\Phi(g)=(g_{S},K,[g_{{\mathcal{C}}}],\mu_{g}),\\ \end{split}$
where $g_{S}$ is the metric on $S$ induced by $g$ and $K$ is the induced second fundamental form (extrinsic curvature) on $S$. The subscript $c$ denotes the compatibility conditions between the initial and boundary data at the corner $\Sigma$.
Due to the diffeomorphism invariance of the vacuum Einstein equations, it is well-known that the map $\Phi$ cannot be one-to-one and so $\Phi$ cannot be an effective parametrization of $\mathbb{E}$. Let $\mathbb{E}^{H}$ be the space of vacuum Einstein metrics in harmonic or wave gauge; cf. §2.3 for details. Let $\mathrm{Diff}_{1}(M)\subset\mathrm{Diff}_{0}(M)$ be the subgroup of diffeomorphisms $\varphi$ such that $\varphi$ is the identity to first order at $S$; equivalently,
$\varphi_{*}(\partial_{t})=\partial_{t}\ \ {\rm on}\ \ S.$
There are no $1^{\rm st}$ order restrictions for $\varphi\in\mathrm{Diff}_{0}(M)$ along ${\mathcal{C}}$. For any metric $g$ such diffeomorphisms leave the future time-like normal vector $\nu_{S}$ to $S$ in $(M,g)$ invariant: $\varphi_{*}(\nu_{S})=\nu_{S}$. It is well-known (cf. again §2.3) that for any $g\in\mathbb{E}$ there is a unique diffeomorphism $\varphi\in\mathrm{Diff}_{1}(M)$ such that
$\varphi^{*}g\in\mathbb{E}^{H}.$
In other words, there is a natural identification
(1.6) ${\mathcal{E}}_{*}=\mathbb{E}/\mathrm{Diff}_{1}(M)\simeq\mathbb{E}^{H},$
where ${\mathcal{E}}_{*}$ is the marked or framed moduli space, consisting of the moduli space ${\mathcal{E}}$ together with the specification of the unit normal $\nu_{S}$ along $S$. The group $\mathrm{Diff}_{0}(M)$ acts on ${\mathcal{E}}_{*}$ with quotient the moduli space ${\mathcal{E}}$,
${\mathcal{E}}_{*}/\mathrm{Diff}_{0}(M)={\mathcal{E}},$
and with fibers given by the orbits of the action of the quotient group $\mathrm{Diff}_{0}(M)/\mathrm{Diff}_{1}(M)$ on ${\mathcal{E}}_{*}$. [^1] It is easy to see that $\mathrm{Diff}_{1}(M)$ is a normal subgroup of $\mathrm{Diff}_{0}(M)$.
The main result of this paper is the following (local-in-time) well-posedness result for the IBVP with boundary data given by (1.3). Let $\mathbb{V}_{S}^{\prime}$ be the space of smooth vector fields along $S$ nowhere tangent to $S$ (the target space for normal vector fields along $S$) and form the product space $\mathcal{I}_{0}\times\mathbb{V}^{\prime}\times{\mathcal{B}}$; let $[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}\subset\mathcal{I}_{0}\times\mathbb{V}^{\prime}\times{\mathcal{B}}$ be the subset consisting of data smoothly compatible at the corner $\Sigma$, cf. §2.4 for details.
Theorem 1.1.
The spaces $\mathbb{E}^{H}$ and $[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$ are smooth tame Frechet manifolds and the map
(1.7) $\begin{split}&\Phi^{H}:\mathbb{E}^{H}\rightarrow[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c},\\ &\Phi^{H}(g)=(g_{S},K,\nu_{S},[g_{{\mathcal{C}}}],\mu_{g}),\\ \end{split}$
is (locally-in-time) a smooth tame diffeomorphism of $\mathbb{E}^{H}$ onto $[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$ with smooth, tame inverse. Thus the IBVP with boundary data (1.3) is well-posed in harmonic gauge.
Briefly, vacuum Einstein metrics on $M$, with arbitrarily given compatible initial and boundary data, exist for a short time, and are uniquely determined in harmonic gauge by their initial data in $\mathcal{I}_{0}\times\mathbb{V}_{S}^{\prime}$ and boundary data in ${\mathcal{B}}$. Further, such solutions depend continuously and in fact smoothly on the target data in $[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$; the IBVP is stable with respect to variation of the target data. A version of Theorem 1.1 holds also off-shell (for non-vacuum Einstein metrics), cf. Theorem 6.1. We refer to [7] for details regarding Frechet spaces and tame maps.
We note here a simple consequence of the proof of Theorem 1.1:
Corollary 1.2.
The marked moduli space ${\mathcal{E}}_{*}$ and the moduli space ${\mathcal{E}}$ of vacuum Einstein metrics on $M$ as above are smooth tame Frechet manifolds.
Note that this result does not depend on any choice of boundary data space ${\mathcal{B}}$; it is intrinsic to the structure of the moduli space ${\mathcal{E}}$, independent of any map to boundary data. The smoothness of the moduli space ${\mathcal{E}}$ has not been established previously. In fact it is known to be false in general for the pure Cauchy problem with compact Cauchy surface $S$ (with $\partial S=\emptyset$); this issue is closely related to the failure of the linearization stability for solutions of the Cauchy problem at Killing initial data, cf. [10].
The boundary data $([g_{{\mathcal{C}}}],\mu_{g})$ of a metric $g$ are natural and simple to compute. However, the boundary data is not fully gauge-invariant or not fully geometric in the sense of [2], [4]. In other words, consider the extension of the map $\Phi$ in (1.5) to general (non-gauged) vacuum Einstein metrics:
(1.8) $\begin{split}&\widehat{\Phi}:\mathbb{E}\rightarrow[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c},\\ &\widehat{\Phi}(g)=(g_{S},K,\nu_{S},[g_{{\mathcal{C}}}],\mu_{g}).\\ \end{split}$
In comparing (1.8) and (1.7), in general,
(1.9) $\widehat{\Phi}\neq\Phi^{H}\circ\pi_{*},$
where $\pi_{*}:\mathbb{E}\rightarrow{\mathcal{E}}_{*}$ is the projection to the orbit space and ${\mathcal{E}}_{*}$ is identified with $\mathbb{E}^{H}$ as in (1.6). The map $\widehat{\Phi}$ is not constant on the fibers of the projection $\pi_{*}$; this is due to the fact that the groups $\mathrm{Diff}_{1}(M)$ and $\mathrm{Diff}_{0}(M)$ act non-trivially on the target data. To explain this in more detail, while the boundary data $[g_{{\mathcal{C}}}]$ is invariant under the action of $\mathrm{Diff}_{1}(M)$ (and $\mathrm{Diff}_{0}(M)$), the volume term $\mu_{g}$ transforms as
(1.10) $(\varphi,\mu_{g})\rightarrow\mu_{\varphi^{*}g}=(\varphi^{*}dv_{g})(dv_{g_{\mathcal{C}}})^{-2/n}.$
For example, in $4$-dimensions, the boundary degrees of freedom for the metric $g$ require $10$ independent boundary conditions, $4$ of which are associated with the choice of gauge. Here, of the remaining $6$ degrees of freedom, $5$ are geometric (corresponding to the specification of $[g_{{\mathcal{C}}}]$), but the scalar volume term $\mu_{g}$ is not (since it is not invariant under the action of $\mathrm{Diff}_{0}(M)$). This may be compared with the works in [6], where only $3$ boundary conditions are geometric (in a restricted sense); the geometric nature of the boundary conditions in [8] is less clear. Moreover, Theorem 1.1 holds in all dimensions while the works in [6] and [8] are restricted to $4$-dimensions.
A related but distinct perspective is to return to the map $\Phi^{H}$ in (1.7) and use the identification (1.6) to obtain via Theorem 1.1 the diffeomorphism
(1.11) $\begin{split}&\Phi_{*}:{\mathcal{E}}_{*}=\mathbb{E}/\mathrm{Diff}_{1}(M)\rightarrow[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c},\\ &\Phi_{*}[g]=(g_{S},K,\nu_{S},[g_{{\mathcal{C}}}],\mu_{g}).\\ \end{split}$
As noted above, the group $\mathrm{Diff}_{0}(M)$ acts on ${\mathcal{E}}_{*}$ with quotient ${\mathcal{E}}={\mathcal{E}}_{*}/\mathrm{Diff}_{0}(M)$ and with fibers given by the orbits of the action of $\mathrm{Diff}_{0}(M)/\mathrm{Diff}_{1}(M)$ on ${\mathcal{E}}_{*}$. The group $\mathrm{Diff}_{0}(M)$ also acts on the target space in the standard way. The map $\Phi_{*}$ is equivariant with respect to these actions and so one obtains an induced map on the quotient spaces:
(1.12) $\begin{split}&\Phi_{*}:{\mathcal{E}}=\mathbb{E}/\mathrm{Diff}_{0}(M)\rightarrow\mathcal{I}_{0}\times_{c}[{\mathcal{B}}],\\ &\Phi_{*}[g]=(g_{S},K,[g_{{\mathcal{C}}}]).\\ \end{split}$
[…]
Note that the boundary data ${\mathcal{B}}$ is invariant under a restricted group of unimodular or volume-preserving diffeomorphisms. Thus, given $g\in\mathbb{E}^{H}$, let
$\mathrm{Diff}_{0}^{dv_{g}}=\{\varphi\in\mathrm{Diff}_{0}(M):\varphi^{*}dv_{g}=dv_{g}\ {\rm at}\ {\mathcal{C}}\},$
be the subgroup of diffeomorphisms in $\mathrm{Diff}_{0}(M)$ which fix the volume form of $g$ at ${\mathcal{C}}$. Let ${\mathcal{O}}_{\mathrm{Diff}_{0}^{dv_{g}}(M)}$ be the orbit of the action of $\mathrm{Diff}_{0}^{dv_{g}}(M)$ on the space of metrics through the metric $g$. We may then form the bundle
$\widetilde{\mathbb{E}}=\cup_{g\in\mathbb{E}^{H}}{\mathcal{O}}_{\mathrm{Diff}_{0}^{dv_{g}}(M)}.$
If $\widetilde{\pi}:\widetilde{\mathbb{E}}\rightarrow\mathbb{E}^{H}$ is the projection map, then one does have
$\Phi|_{\widetilde{\mathbb{E}}}=\Phi^{H}\circ\widetilde{\pi}:\widetilde{\mathbb{E}}\rightarrow\mathcal{I}_{0}\times_{c}{\mathcal{B}}.$
Thus $\Phi|_{\widetilde{\mathbb{E}}}$ is constant on the fibers of $\widetilde{\pi}$ and is a diffeomorphism transverse to the fibers by Theorem 1.1. However, this does not hold on the full space $\mathbb{E}$, as noted in (1.9).
In general Theorem 1.1 implies two vacuum Einstein metrics $g_{1},g_{2}\in\mathbb{E}$ are isometric (within the gauge group $\mathrm{Diff}_{0}(M)$) if and only if there are diffeomorphisms $\varphi_{i}\in\mathrm{Diff}_{0}(M)$, $i=1,2$, such that $\varphi_{i}^{*}g_{i}\in\mathbb{E}^{H}$ and
(1.13) $\Phi^{H}(\varphi_{1}^{*}g_{1})=\Phi^{H}(\varphi_{2}^{*}g_{2}).$
This requires that $g_{1}$ and $g_{2}$ have the same initial data, induce the same conformal structure on the boundary ${\mathcal{C}}$ and
$\varphi_{1}^{*}\mu_{g_{1}}=\varphi_{2}^{*}\mu_{g_{2}},$
i.e.
$\varphi_{1}^{*}dv_{g_{1}}(dv_{(g_{1})_{{\mathcal{C}}}})^{-2/n}=\varphi_{2}^{*}dv_{g_{2}}(dv_{(g_{2})_{{\mathcal{C}}}})^{-2/n}.$
We point out that the uniqueness for the gauged map $\Phi^{H}$ in (1.13) cannot be improved to the ungauged map $\Phi$ in (1.5). Thus it is not true that for general $g_{1},g_{2}\in\mathbb{E}$ (not necessarily in harmonic gauge),
(1.14) $\Phi(g_{2})=\Phi(g_{1})\Rightarrow g_{2}=\varphi^{*}g_{1},$
for some $\varphi\in\mathrm{Diff}_{0}(M)$, i.e. $\Phi$ in (1.5) cannot be one-to-one modulo isometry in $\mathrm{Diff}_{0}(M)$. To see this, suppose $g^{1},g^{2}$ are a pair of vacuum Einstein metrics with the same initial data and the same Dirichlet boundary data so that $[g^{1}_{{\mathcal{C}}}]=[g^{2}_{{\mathcal{C}}}]$ and $dv_{g^{1}_{{\mathcal{C}}}}=dv_{g^{2}_{{\mathcal{C}}}}$. As noted above, $\mathrm{Diff}_{0}(M)$ acts transitively on the space of bulk volume forms $dv_{g}$ restricted to ${\mathcal{C}}$. Hence, there exists $\psi\in\mathrm{Diff}_{0}(M)$ such that for $\widetilde{g}=\psi^{*}g^{2}$, $dv_{\widetilde{g}}=\psi^{*}dv_{g^{2}}=dv_{g^{1}}$. Since $\mathrm{Diff}_{0}(M)$ leaves any volume form $dv_{g_{{\mathcal{C}}}}$ invariant, it follows that
$\Phi(\widetilde{g})=\Phi(g^{1}).$
Hence the validity of (1.14) would imply $\widetilde{g}=\varphi^{*}g^{1}$ and so $g^{2}=(\varphi\circ\psi^{-1})^{*}g^{1}$. Thus (1.14) implies uniqueness of solutions to the Dirichlet boundary value problem in ${\mathcal{E}}$. However, this was proved to be false in general, (i.e. without further conditions), in [3], [4]; it follows that (1.14) is also false in general.
The proof of Theorem 1.1 follows the general lines of the main results in [2, 3]. In particular, after a detailed analysis of the linearized problem, the non-linear result follows from the Nash-Moser inverse function theorem.
A brief summary of the contents of the paper is as follows. In §2, we introduce preliminary material needed for the work to follow. Much of the discussion in this section follows [2], [3] closely, but due to the use of the Nash-Moser theorem is nevertheless quite extensive. In §3, we derive the basic equations for the boundary data not part of the Dirichlet boundary data and prove the local existence theorem (Theorem 3.2). The important apriori energy estimates are derived in §4. The work in §3 and §4 is then extended to the global setting of a general compact Cauchy surface in §5. The main results are then proved in §6 via the Nash-Moser theorem.
2. Initial Material
The overall approach to the proof of Theorem 1.1 is similar to the approaches taken in Parts I and II, [2, 3]. Although this section makes this work self-contained, we refer to [2, 3] for some further details and proofs.
2.1. Basic definitions and notation.
Let $S$ be a compact, connected and oriented $n$-manifold with boundary $\partial S=\Sigma$; here the boundary $\Sigma$ may or may not be connected. Let $M=I\times S$, where $I=[0,1]$ is the range of a time function $t:M\rightarrow I$. Let $S_{t}=\{t\}\times S$; $S_{t}$ is identified with $S_{0}$ upon introduction of local coordinates. We consider smooth Lorentz metrics $g$ on $M$ which are globally hyperbolic in the sense of manifolds with boundary, where the boundary ${\mathcal{C}}=I\times\Sigma$ is timelike with respect to the metric $g$; the induced metric $g_{{\mathcal{C}}}$ is thus a globally hyperbolic metric on ${\mathcal{C}}$ with closed Cauchy surface $\Sigma$.
Let $\nu_{S}$ denote the future timelike unit normal vector field to $S\subset(M,g)$ and similarly let $\nu_{{\mathcal{C}}}$ denote the (spacelike) outward unit normal to ${\mathcal{C}}\subset(M,g)$. Let
$K=\tfrac{1}{2}{\mathcal{L}}_{\nu_{S}}g|_{S}\ {\rm and}\ A=\tfrac{1}{2}{\mathcal{L}}_{\nu_{{\mathcal{C}}}}g|_{{\mathcal{C}}},$
denote the extrinsic curvatures (second fundamental forms) of $S$ and ${\mathcal{C}}$ in $(M,g)$ respectively.
In local coordinates $x^{\alpha}=(t,x^{a})$ near the corner $\Sigma$, $x^{0}=t\geq 0$ is a defining function for $S=\{t=0\}$, while $x^{1}\leq 0$ is chosen to be a defining function for ${\mathcal{C}}=\{x^{1}=0\}$. The coordinates $x^{\alpha},\alpha=2,\cdots,n$ are local coordinates of the corner $\Sigma$ itself. As usual, Greek letters $\alpha,\beta$ denote spacetime indices $0,\dots,n$. Unless otherwise indicated, Roman letters denote spacelike indices $1,\dots,n$ and capital Roman letters denote corner indices $2,\dots,n$.
We will use the same function spaces as in [3], namely the standard Sobolev spaces. Let $D$ denote a domain contained either in $M$, $S$, ${\mathcal{C}}$, $\Sigma$ or the corresponding $t$-level sets $S_{t}$, $\Sigma_{t}$, or the corresponding $t$-sublevel sets, $M_{t}$, ${\mathcal{C}}_{t}$; $C_{t}=\{p\in{\mathcal{C}}:t(p)\leq t\}$ and similarly for $M_{t}$. Define the Sobolev $H^{s}$ norm on functions on $D$ by
[…]
where $\partial_{D}^{k}$ consists of all partial coordinate derivatives tangent to $D$ of order $\leq s$. In place of coordinate derivatives, one may use the components of covariant derivatives of $v$ up to order $s$, given a (background) Riemannian metric $g_{D}$ on $D$. The volume form $dv_{D}$ is also induced from the metric $g_{D}$. We will also use the generalizations of the spaces $H^{s}$ for real-valued $s\in\mathbb{R}^{+}$.
Define the stronger $\bar{H}^{s}$ norm by including all space-time derivatives up to order $s$, so that
[…]
where $\partial_{M}$ denotes partial derivatives along all coordinates of $M$ at $D$.
Finally, for the Cauchy slices $S_{t}$, define the boundary stable $H^{s}$ norm on $S_{t}$ by
$||v||_{\bar{\mathcal{H}}^{s}(S_{t})}^{2}=||v||_{\bar{H}^{s}(S_{t})}^{2}+||v||_{\bar{H}^{s}({\mathcal{C}}_{t})}^{2}.$
We also recall the Sobolev trace theorem for the restriction map $f:D\rightarrow\mathbb{R}$ to $f|_{\partial D}:\partial D\rightarrow\mathbb{R}$:
$||f||_{H^{s-1/2}(\partial D)}\leq C||f||_{H^{s}(D)}.$
It will always be assumed that $s$ is sufficiently large, depending on $n$, so that by the Sobolev embedding theorem, $H^{s}\subset C^{2}$.
We will use the function space
(2.1) ${\mathcal{N}}^{s}(M)=C^{0}(I,\bar{\mathcal{H}}^{s}(S)),$
for functions on $M=M_{1}$. With the usual $C^{0}$ norm in $I$, the space ${\mathcal{N}}^{s}$, with associated norm $||\cdot||_{{\mathcal{N}}^{s}}$ is a separable Banach space. The normed topology (2.1) is commonly used for the space of solutions of wave-type equations.
Note that
$H^{s+1}(M)\subset{\mathcal{N}}^{s}(M)\subset H^{s}(M).$
Let $Met^{s}(M)$ be the space of globally hyperbolic Lorentz metrics $g$ on $M$, with coefficients $g_{\alpha\beta}\in{\mathcal{N}}^{s}$ in a smooth atlas for $M$. Similarly, let $(S^{2}(M))^{s-2}$ denote the space of symmetric bilinear forms on $M$ with coefficients in $H^{s-2}(M)$. The initial data of $g$ is given by the pair $(g_{S},K)$ consisting the metric $g_{S}$ induced on the initial Cauchy surface $S=S_{0}$ and the extrinsic curvature $K$ of $(S,g_{S})\subset(M,g)$. This pair takes its values in the space $\mathcal{I}^{s}=Met^{s}(S)\times(S^{2}(S))^{s-1}$ of all initial data $(\gamma,\kappa)$, with the $H^{s}(S)\times H^{s-1}(S)$ topology.
For this work, the space of boundary data is
${\mathcal{B}}^{s-1/2}={\rm Conf}^{s-1/2}({\mathcal{C}})\times H_{+}^{s-1/2}({\mathcal{C}}).$
The first factor consists of the space of pointwise conformal classes of globally hyperbolic Lorentz metrics on the boundary ${\mathcal{C}}$, with coefficients in $H^{s-1/2}({\mathcal{C}})$. The second factor denotes positive $H^{s-1/2}$ functions on ${\mathcal{C}}$. The bulk metric $g$ induces a boundary metric $g_{{\mathcal{C}}}$ and scalar $\mu_{g}=dv_{g}(dv_{g_{{\mathcal{C}}}})^{-2/n}$ with
$([g_{{\mathcal{C}}}],\mu_{g})\in{\mathcal{B}}^{s-1/2}.$
Here $\mu_{g}$ is identified with a scalar function on ${\mathcal{C}}$ by the choice of a fixed but otherwise arbitrary smooth volume form on $M$ and an induced volume form on the boundary ${\mathcal{C}}$.
We consider the map
(2.2) $\Phi:Met(M)\rightarrow[S^{2}(M)\times\mathcal{I}\times{\mathcal{B}}]_{c}:={\mathcal{T}},$
$\Phi(g)=(\mathrm{Ric}_{g},g_{S},K,[g_{{\mathcal{C}}}],\mu_{g}),$
where $(g_{S},K)$ and $([g_{{\mathcal{C}}}],\mu_{g})$ are the initial and ${\mathcal{B}}$-valued boundary data induced by $g$. The subscript $c$ denotes the compatibility conditions between the initial and boundary data at the corner $\Sigma$; these are discussed in detail in §2.3-2.4. For simplicity, we have dropped here the notation for the topologies, indexed by $s$.
This is the extension of the map $\Phi$ in (1.5) to the space of all globally hyperbolic Lorentz metrics on $M$.
2.2. Localization
As usual with hyperbolic PDE problems, the main existence results will first be proved locally, in regions where the constant coefficient approximation is effective. Such local solutions are then patched together, by a standard and well-known process, via rescaling and a partition of unity,
Given a metric $g$ on $M$ as in §1, the localization at a point $p\in\Sigma$ is a (small) neighborhood $U\subset M$ of $p$, diffeomorphic via a local chart to a Minkowski corner
$\mathbf{R}=\{(t=x^{0},x^{1},\dots,x^{n}):t\geq 0,x^{1}\leq 0\},$
with $S\cap U\subset\{t=0\}$, ${\mathcal{C}}\cap U\subset\{x^{1}=0\}$ and $x^{\alpha}(p)=0$.
In such a metrically small neighborhood $U$, the metric and coordinates are renormalized simultaneously by rescaling; thus for $\lambda$ small, set
$\widetilde{g}=\lambda^{-2}g,\ \ \widetilde{x}^{\alpha}=\lambda^{-1}x^{\alpha}$
so that $\partial_{\widetilde{x}^{\alpha}}=\lambda\partial_{x^{\alpha}}$ and
(2.3) $\widetilde{g}(\partial_{\widetilde{x}^{\alpha}},\partial_{\widetilde{x}^{\beta}})|_{\widetilde{x}}=g(\partial_{x^{\alpha}},\partial_{x^{\beta}})|_{\lambda x},$
and similarly for variations $h$, $\widetilde{h}$ of $g$ and $\widetilde{g}$. While the components $g_{\alpha\beta}$ of $g$ are invariant under such a rescaling, all higher derivatives become small:
(2.4) $\partial_{\widetilde{x}^{\mu}}^{k}\widetilde{g}_{\alpha\beta}=\lambda^{k}\partial_{x^{\mu}}^{k}g_{\alpha\beta}=O(\lambda^{k}).$
Thus (for smooth metrics $g$) the coefficients are smoothly close to constant functions in the rescaled chart,
(2.5) $||\widetilde{g}-g_{\alpha_{0}}||_{C^{\infty}(U)}\leq\varepsilon=\varepsilon(\lambda,g),$
where $g_{\alpha_{0}}$ is a flat (constant coefficent) Minkowski metric. This is the frozen coefficient approximation.
The $\lambda\rightarrow 0$ limit blow-up metric $g_{\alpha_{0}}$ is the Minkowski metric, given in adapted coordinates as above by
(2.6) $g_{\alpha_{0}}=-dt^{2}-\alpha_{0}dtdx^{1}+\sum_{i=1}^{n}(dx^{i})^{2}.$
Here $\alpha_{0}=\alpha(p)=\langle\nu_{S},\nu_{{\mathcal{C}}}\rangle_{p}$ is the value of the (scale-invariant) corner angle at $p$. Note that for $g_{\alpha_{0}}$,
(2.7) $\nu_{S}=-\nabla t=(1+\alpha_{0}^{2})^{-1/2}(\partial_{0}+\alpha_{0}\partial_{1}),\ \ \nu_{\mathcal{C}}=\nabla x^{1}=(1+\alpha_{0}^{2})^{-1/2}(\partial_{1}-\alpha_{0}\partial_{0}).$
Note also that the leading order Sobolev norms scale as follows: for $D\subset S$ or $D\subset{\mathcal{C}}$ of dimension $n$,
$||f||_{H_{\widetilde{g}}^{s}(D)}^{2}=\lambda^{2s-n}||f||_{H_{g}^{s}(D)}^{2},$
$||f||_{{\mathcal{N}}_{\widetilde{g}}^{s}(M)}^{2}=\lambda^{2s-n}||f||_{{\mathcal{N}}_{g}^{s}(M)}^{2}.$
Also as in [2], $U$ is always assumed to be smoothly embedded in a larger region $\widetilde{U}$,
$U\subset\widetilde{U},$
with $\widetilde{U}$ still covered by the adapted coordinates $(t,x^{i})$, with $t\geq 0$ and $x^{1}\leq 0$ in $\widetilde{U}$ so that the initial surface $S$, boundary ${\mathcal{C}}$ and corner $\Sigma$ in $\widetilde{U}$ are an extension of the corresponding domains in $U$. We also assume (2.5) still holds in $\widetilde{U}$. All target data in ${\mathcal{T}}$ as in (2.2) and later in ${\mathcal{T}}^{H}$ which is given in $U$ is extended off $U$ to be of compact support in $\widetilde{U}$ away from $S\cap U$ and ${\mathcal{C}}\cap U$. Thus in such localizations, all target data vanishes in a neighborhood of the full timelike boundary of $\widetilde{U}$ and in a neighborhood of the initial slice $\{t=0\}$ away from ${\mathcal{C}}\cap U$ and $S\cap U$ respectively. The same statements hold for variations of the target data, i.e. in the tangent spaces $T({\mathcal{T}})$ or $T({\mathcal{T}}^{H})$.
For later reference, we note that solutions $h$ of linear systems of wave equations on $\widetilde{U}$ appearing below and throughout §3 - §4 enjoy the finite propagation speed property; if the initial and boundary data of $h$ are compactly supported in $\widetilde{U}$, then $h$ is also compactly supported in $\widetilde{U}$, for some definite (possibly small) time $t>0$.
2.3. Gauge analysis.
Due to gauge issues, the map $\Phi$ in (2.2) is well-known to be neither surjective nor injective, or given by non-linear hyperbolic evolution equations. A choice of gauge is used to deal with these issues. This will be described here by a choice of vector field $V$ on $M$ and replacing the vacuum Einstein equations (1.1) by the reduced Einstein equations
(2.8) $\mathrm{Ric}_{g}+\delta^{*}V=0,$
where $\delta^{*}V=\frac{1}{2}{\mathcal{L}}_{V}g$. As in [2, 3] we choose here the (generalized) harmonic or wave coordinate gauge, given by
(2.9) $V_{g}=\Box_{g}x^{\alpha}\partial_{x^{\alpha}},$
where $\Box_{g}$ is the d’Alembertian or wave operator with respect to $g$. It is well-known that the gauge (2.9) can be naturally globalized to all of $M$ by choice of a wave map to a fixed target space, cf. [2] for instance.
The gauged version of the map $\Phi$ in (2.2) is then given by
(2.10) $\Phi^{H}:Met(M)\rightarrow{\mathcal{T}}^{H},$
$\Phi^{H}(g)=\big{(}(\mathrm{Ric}_{g}+\delta_{g}^{*}V_{g}),(g_{S},K,\nu_{S},V|_{S}),([g_{{\mathcal{C}}}],\mu_{g},V|_{{\mathcal{C}}})\big{)},$
where $V|_{S}$ and $V|_{{\mathcal{C}}}$ are the restrictions of $V_{g}$ to $S$ and ${\mathcal{C}}$ respectively and $\nu_{S}$ is the future unit normal to $S$ in $(M,g)$. In $4$-dimensions, the initial data $(g_{S},K,\nu_{S},V|_{S})$ consists of the (required) 20 degrees of freedom, while the boundary conditions $([g_{{\mathcal{C}}}],\mu_{g},V|_{{\mathcal{C}}})$ consist of the (required) 10 degrees of freedom. Similar consistency holds in all dimensions.
The target space ${\mathcal{T}}^{H}$ in (2.10) is given as follows. Let ${\mathcal{U}}$ denote the uncoupled space of target data, so that
${\mathcal{U}}=S^{2}(M)\times\mathcal{I}\times\mathbb{V}_{S}^{\prime}\times\mathbb{V}_{S}\times{\mathcal{B}}\times\mathbb{V}_{{\mathcal{C}}}.$
This is a product space, extending the product space in (2.2); $\mathbb{V}^{\prime}_{S}$ is the space of $H^{s}$ vector fields along $S$ nowhere tangent to $S$, $\mathbb{V}_{S}$ is the space of $H^{s-1}$ vector fields $V_{S}$ along $S$ and $\mathbb{V}_{{\mathcal{C}}}$ is the space of $H^{s-3/2}$ vector fields $V_{{\mathcal{C}}}$ along ${\mathcal{C}}$. By construction, ${\mathcal{U}}$ is a separable Hilbert space, and Frechet space when $s=\infty$, i.e. for $C^{\infty}$ data.
The metric $g$ induces data in ${\mathcal{U}}$ through the map $\Phi^{H}$. The two equations for $g$, namely the bulk equation $\mathrm{Ric}_{g}+\delta^{*}V_{g}=Q$ and the definition of the gauge field $V_{g}$ in (2.9), as well as definition of the normal vector $\nu_{S}$, determine compatibility conditions for the data in ${\mathcal{U}}$ at the corner $\Sigma$. These compatibility conditions are expressed in terms of equations on the data in ${\mathcal{U}}$. The subspace of compatible target data, given by
${\mathcal{U}}_{c}=[S^{2}(M)\times\mathcal{I}\times\mathbb{V}_{S}^{\prime}\times\mathbb{V}_{S}\times{\mathcal{B}}\times\mathbb{V}_{{\mathcal{C}}}]_{c}\subset{\mathcal{U}},$
is then the subspace satisfying these compatibility conditions. The compatibility conditions are expressed as the zero-locus of a (large) collection of functions on ${\mathcal{U}}$, ordered according to the degree of differentiability.
Note that the initial gauge velocity $\partial_{t}V|_{S}$ is not prescribed by the data in ${\mathcal{U}}$ or ${\mathcal{U}}_{c}$. In fact, the gauge velocity at $S$ is only prescribed by imposing the constraint equations, i.e. the Gauss and Gauss-Codazzi equations, on $S$:
(2.11) $|K|^{2}-(\mathrm{tr}_{g_{S}}K)^{2}-R_{g_{S}}=R_{g}-2\mathrm{Ric}_{g}(\nu_{S},\nu_{S})=\mathrm{tr}_{g_{S}}\mathrm{Ric}_{g}-3\mathrm{Ric}_{g}(\nu_{S},\nu_{S}),$
(2.12) ${\rm div}_{g_{S}}(K-Hg_{S})=\mathrm{Ric}_{g}(\nu_{S},\cdot).$
To express this in terms of target data in ${\mathcal{U}}_{c}$, denote
$Q=\mathrm{Ric}+\delta^{*}V,$
so that (2.11)-(2.12) becomes
(2.13) $|K|^{2}-(\mathrm{tr}_{g_{S}}K)^{2}-R_{g_{S}}=\mathrm{tr}_{g_{S}}(Q-\delta^{*}V)-3(Q(\nu_{S},\nu_{S})-\delta^{*}V(\nu_{S},\nu_{S})),$
(2.14) ${\rm div}_{g_{S}}(K-Hg_{S})=Q(\nu_{S},\cdot)-\delta^{*}V(\nu_{S},\cdot).$
It is straightforward to check that $\delta^{*}V$ is determined along $S$ by the initial data $(V|_{S},\partial_{t}V|_{S})$. Let then $\mathbb{V}_{S}$, the space of vector fields along $S$, representing the target space for $\partial_{t}V|_{S}$ and consider the constraint operator
$C:{\mathcal{U}}_{c}\times\mathbb{V}_{S}\rightarrow\mathbb{V}_{S},$
(2.15) $\begin{split}&C(Q,\gamma_{S},\kappa,V_{S},\partial_{t}V_{S},\nu_{S},\gamma_{{\mathcal{C}}},V_{{\mathcal{C}}})=\begin{cases}|\kappa|^{2}-(\mathrm{tr}_{\gamma_{S}}\kappa)^{2}-R_{\gamma_{S}}-\mathrm{tr}_{g_{S}}(Q-\delta^{*}V)+3(Q(\nu_{S},\nu_{S})-\delta^{*}V(\nu_{S},\nu_{S})),\\ {\rm div_{\gamma_{S}}}\kappa-d_{S}(\mathrm{tr}_{\gamma_{S}}\kappa)-Q(\nu_{S},\cdot)+\delta^{*}V(\nu_{S},\cdot).\\ \end{cases}\end{split}$
We then define
(2.16) ${\mathcal{T}}^{H}=({\mathcal{U}}_{c}\times\mathbb{V}_{S})_{0},$
to be the zero-set of the constraint operator $C$. By inspection in (2.15), on ${\mathcal{T}}^{H}$, $\partial_{t}V_{S}$ is uniquely determined by the remaining data in ${\mathcal{U}}_{c}$, and so ${\mathcal{T}}^{H}$ is naturally identified with ${\mathcal{U}}_{c}$.
The overall method of proof of well-posedness is to prove that $\Phi^{H}$ in (2.10) is a tame Fredholm map with a tame inverse at the linearized level and to then apply the Nash inverse function theorem in Frechet spaces. Thus, most of the analysis to follow deals with the study of the linearization
(2.17) $D\Phi_{g}^{H}:T_{g}Met(M)\rightarrow T_{\tau}({\mathcal{T}}^{H}),$
where $\tau=\Phi^{H}(g)$.
The bulk term in the linearization $D\Phi_{g}^{H}$ is given by
(2.18) $L(h)=\mathrm{Ric}^{\prime}_{h}+\delta^{*}V^{\prime}_{h}+(\delta^{*})^{\prime}_{h}V.$
The linearization of the gauge field $V=V_{g}$ is given by
(2.19) $V^{\prime}_{h}=\beta_{g}h-\langle D^{2}x^{\alpha},h\rangle\partial_{x^{\alpha}},$
where $\beta_{g}=\delta_{g}+\frac{1}{2}d\mathrm{tr}_{g}h$ is the Bianchi operator. The bulk linearized equation $L(h)=F$ then has the form
(2.20) $L(h)=\tfrac{1}{2}D^{*}Dh+P(h)=F,$
where $P$ is a first order linear differential operator on $h$ and $D^{*}D=-\nabla_{e_{\alpha}}\nabla_{e_{\alpha}}+\nabla_{\nabla_{e_{\alpha}}e_{\alpha}}$ is the tensorial wave operator on symmetric bilinear forms. Of course all the coefficients of the derivatives in (2.20) depend on the background metric $g$ at which the linearization is formed.
The equation (2.20) may be written in local coordinates in the form
(2.21) $(L(h))_{\alpha\beta}=-\tfrac{1}{2}\Box_{g}(h_{\alpha\beta})+P_{\alpha\beta}(h)=F_{\alpha\beta}.$
where $P(h)$ involves $h$ and its first derivatives. In the localized context discussed in §2.2, the coefficients of $P$ are all of order $\varepsilon$, $P(h)=\varepsilon(\partial h,h)$. Here $\Box_{g}f=\frac{1}{\sqrt{-g}}\partial_{\alpha}(g^{\alpha\beta}\sqrt{-g}\partial_{\beta}f)$ is the D’Alembertian or wave operator with respect to $g$.
In later sections, we will repeatedly need to solve linear wave equations of the form (2.21), i.e.
(2.22) $-\tfrac{1}{2}\Box_{g}v+p(v)=\varphi,$
for given $\varphi$. Here $v$ may be either scalar valued or vector-valued and again $p(v)$ is first order in $v$. Although solvability of the linear equation or system (2.22) holds more generally, we will only need to actually solve such systems in $C^{\infty}$, so with $g$, $p$ and $\varphi$ in $C^{\infty}$.
The next Lemma is well-known and will be used repeatedly in §3 and §4.
Lemma 2.1.
For Dirichlet boundary data, the $C^{\infty}$ system (2.22) is solvable for any smooth $\varphi$, with standard energy estimates.
Proof.
This is a standard and well-known result, cf. [5], [11] for further details. Of course it is assumed here implicitly that the initial data and boundary data for $v$ satisfy the necessary smooth compatibility conditions (as above) at the corner $\Sigma$, cf. also §2.4. Energy estimates for such equations are also standard; for later reference, we provide the main result here. Thus, the strong or boundary stable energy estimate with Dirichlet boundary data states that any (smooth) solution of (2.22) satisfies the bound
(2.23) $||v||_{\bar{\mathcal{H}}^{s}(S_{t})}^{2}\leq C[||v||_{H^{s}(S_{0})}^{2}+||\partial_{t}v||_{H^{s-1}(S_{0})}^{2}+||v||_{H^{s}({\mathcal{C}}_{t})}^{2}+||\varphi||_{H^{s-1}(M_{t})}^{2}],$
where $t\in[0,1]$ and $C$ is a constant depending only on $g$ and the coefficients of $p$. Note the difference in the stronger and weaker norms on the left and right of (2.23). We recall that ${\mathcal{C}}_{t}=\{p\in{\mathcal{C}}:t(p)\leq t\}$ and similarly for $M_{t}$.
Remark 2.2.
One also has existence and uniqueness for smooth solutions of the equation (2.22) with given Neumann boundary data $b(v)=\nu_{{\mathcal{C}}}(v)$. However, in this case, there is no effective energy estimate as in (2.23); instead the energy estimate only holds with a loss of a (fractional) derivative. The optimal result in general is:
(2.24) $|v||_{\bar{\mathcal{H}}^{s}(S_{t})}^{2}\leq C[||v||_{H^{s}(S_{0})}^{2}+||\partial_{t}v||_{H^{s-1}(S_{0})}^{2}+||\nu_{{\mathcal{C}}}(v)||_{H^{s-1/3}({\mathcal{C}}_{t})}^{2}+||\varphi||_{H^{s-1}(M_{t})}^{2}],$
so there is a ’loss’ of $2/3$ of a derivative. We refer to [9, Ch.8], and in particular to [12, Thm.9] for the proof and further details.
We next discuss the gauge term $V$ and its linearization $V^{\prime}_{h}$. Consider the bulk equation
(2.25) $\mathrm{Ric}_{g}+\delta_{g}^{*}V=Q,$
where $Q$ is given data in the target data space ${\mathcal{T}}^{H}$. We will need the following results on the gauge field from [2]. For completeness we state the full result but refer to [2, [Lemmas 3.2-3.3] for the proof.
Lemma 2.3.
The gauge field $V$ satisfies the equation
(2.26) $-\tfrac{1}{2}(\Box V+\mathrm{Ric}_{g}(V))=\beta(Q).$
on $M$, where $\Box$ denotes the vector wave operator $\Box=-D^{*}D$ acting on vector fields. The initial data $V|_{S}$, $\partial_{t}V|_{S}$ and boundary data $V|_{{\mathcal{C}}}$ are determined by the target data in ${\mathcal{T}}^{H}$. Thus such data uniquely determine $V$.
Similarly, for the linearization $L(h)=F$ as in (2.20), one has
(2.27) $-\tfrac{1}{2}(\Box_{g}V^{\prime}_{h}+\mathrm{Ric}_{g}(V^{\prime}_{h}))+\tfrac{1}{2}\nabla_{V}V^{\prime}_{h}=\beta(F)+\beta^{\prime}_{h}\mathrm{Ric}+O_{2,1}(V,h),$
where $O_{2,1}(V,h)$ is second order in $V$, first order in $h$ and vanishes if $V=0$.
The initial data $V^{\prime}_{h}|_{S}$, $\partial_{t}V^{\prime}_{h}|_{S}$ and boundary data $V^{\prime}_{h}|_{{\mathcal{C}}}$ are determined by the target space data in $T({\mathcal{T}}^{H})$.
As in [2], decompose the gauge variation $V^{\prime}_{h}$ as a sum
(2.28) $V^{\prime}_{h}=V^{\prime}_{F}+W^{\prime}_{h},$
where $V^{\prime}_{F}=V^{\prime}_{F}({\mathcal{T}})$ is the vector field satisfying
(2.29) $-\tfrac{1}{2}[\Box V^{\prime}_{F}+\mathrm{Ric}(V^{\prime}_{F})]+\tfrac{1}{2}\nabla_{V}V^{\prime}_{F}=\beta(F),$
with $(V^{\prime}_{F})|_{S}=(V^{\prime}_{h})|_{S}$, $(\partial_{t}V^{\prime}_{F})|_{S}=(\partial_{t}V^{\prime}_{h})|_{S}$ along $S$ and $(V^{\prime}_{F})|_{{\mathcal{C}}}=V^{\prime}_{h}|_{{\mathcal{C}}}$ along ${\mathcal{C}}$. Thus $V^{\prime}_{F}$ is uniquely determined by the target space data $T({\mathcal{T}}^{H})$ (so is independent of $h$). The remaining component $W^{\prime}_{h}$ then satisfies
(2.30) $-\tfrac{1}{2}[\Box W^{\prime}_{h}+\mathrm{Ric}(W^{\prime}_{h})]+\tfrac{1}{2}\nabla_{V}W^{\prime}_{h}=\beta^{\prime}_{h}\mathrm{Ric}+O_{2,1}(V,h),$
with $W^{\prime}_{h}|_{S}=\partial_{t}W^{\prime}_{h}|_{S}=W^{\prime}_{h}|_{{\mathcal{C}}}=0$.
Remark 2.4.
(i). Lemma 2.3 implies the standard result that solutions of the gauge-reduced Einstein equations
$\mathrm{Ric}_{g}+\delta^{*}V=0,$
with target data $V=0$ at $S\cup{\mathcal{C}}$ and with $g$ satisfying the vacuum constraint equations (2.11)-(2.12) along $S$ necessarily satisfy $V=0$ on $M$, and so are solutions of the Einstein equations
$\mathrm{Ric}_{g}=0$
on $M$. The same result holds locally, in the context of the localization in §2.2, as well as for the linearized equations (2.19).
Conversely, given any vacuum Einstein metric $g$, there is a unique diffeomorphism $\varphi\in\mathrm{Diff}_{1}(M)$ such that $\widetilde{g}=\varphi^{*}g$ is in harmonic gauge, $V_{\widetilde{g}}=0$. Recall that $\mathrm{Diff}_{1}(M)\subset\mathrm{Diff}_{0}(M)$ is the group of smooth diffeomorphisms equal to the identity to first order on $S$.
(ii). The initial data $g_{\alpha\beta}$, $\partial_{t}g_{\alpha\beta}$ for $g_{\alpha\beta}$ along $S$ are determined by the initial data $g_{S}$, $K$, $\nu_{S}$ and $V|_{S}$ along $S$, and similarly for linearizations $h$, cf. [2].
(iii). For later use, note that the gauge variation $V^{\prime}_{h}$ satisfies the boundary stable energy estimate (2.23), since the Dirichlet boundary value $V^{\prime}_{h}|_{{\mathcal{C}}}$ is part of the target data in $T({\mathcal{T}}^{H})$.
2.4. Geometry at the corner
As noted in [2], the conformal factor $\varphi$ of the given boundary data $[g_{{\mathcal{C}}}]$ is determined at the corner $\Sigma$ by the initial data $g_{S}$. Consequently, the full metric $g_{\mu\nu}$ at $\Sigma$ is determined at the corner except possibly for the corner angle
(2.31) $\alpha=\langle\nu_{S},\nu_{{\mathcal{C}}}\rangle.$
In fact for the boundary data $([g_{{\mathcal{C}}}],H_{{\mathcal{C}}})\in{\mathcal{B}}_{H}$, the angle $\alpha$ is not determined by initial and boundary data, cf. again [2]. On the other hand, it was shown in [3] that $\alpha$ is determined by initial and boundary data when the boundary data ${\mathcal{B}}={\mathcal{B}}_{Dir}$ is Dirichlet boundary data. The same occurs here.
Lemma 2.5.
In local coordinates $\{x^{\mu}\}$ near $\Sigma$ with $g_{11}=f^{2}$ and $g_{1a}=0$, $a=2,\dots,n$ (zero shift), one has
(2.32) $dv_{g}=\sqrt{f^{2}det(-g_{{\mathcal{C}}})+\alpha^{2}det(g_{\Sigma})}dx^{0}\wedge\cdots\wedge x^{n},$
where $\alpha=-g_{01}$. Consequently,
(2.33) $\mu_{g}=\frac{\sqrt{f^{2}det(-g_{{\mathcal{C}}})+\alpha^{2}det(g_{\Sigma})}}{[det(-g_{{\mathcal{C}}})]^{1/n}},$
and so at the corner $\Sigma$, $\alpha$ is uniquely determined, up to sign, by the initial and boundary data in $\mathcal{I}\times{\mathcal{B}}$.
Proof.
The formula (2.32) is a simple calculation in local coordinates, as is the relation $\alpha=-g_{01}$. As noted above, at $\Sigma$, both $f^{2}=g_{11}$ and the conformal factor $\varphi$ are determined by initial data. Hence, $g_{{\mathcal{C}}}$ and $g_{\Sigma}$ are determined at $\Sigma$ by the target data in $\mathcal{I}\times{\mathcal{B}}$. Thus the second statement follows from the formula (2.33). ∎
Remark 2.6.
A change $\alpha\rightarrow-\alpha$ of the sign of $\alpha$ corresponds to a change in time orientation $\partial_{t}\rightarrow-\partial_{t}$ and so this distinction will be ignored in the following.
Remark 2.7.
Let $\sigma$ denote a fixed background metric in the conformal class $[g_{{\mathcal{C}}}]$, so that $g_{{\mathcal{C}}}=\varphi^{2}\sigma$, where $[\sigma]$ is determined by the target data but $\varphi$ is not. As noted above, $\varphi|_{\Sigma}$ is determined by the initial and boundary data. Similarly, the initial velocity of $\varphi$ at $\Sigma$ is determined by the target data; in fact from [2], one has
(2.34) $(n-1)T(\varphi)=\sqrt{1+\alpha^{2}}\mathrm{tr}_{\Sigma}K_{S}-\alpha H_{\Sigma}-H_{\sigma}.$
where $T$ is the unit timelike normal to $\Sigma$ in $({\mathcal{C}},g_{{\mathcal{C}}})$ and $H_{\sigma}$ is the mean curvature of $\Sigma$ in $({\mathcal{C}},\sigma)$. Note that all the terms on the right side are determined by the initial and boundary data. The linearization of this result is used in Lemma 3.1 below.
The compatibility conditions on the initial and boundary data at the corner $\Sigma$ were analysed in detail in [3] for Dirichlet boundary data; the analysis for the conformal-volume or twisted Dirichlet boundary data $([g_{{\mathcal{C}}}],\mu_{g})$ proceeds in exactly the same way with only minor changes and so we refer to [3] details. In particular we note the following result, but refer to [3] for the proof.
Proposition 2.8.
With $C^{\infty}$ data, the target space ${\mathcal{T}}^{H}$ in (2.16) is a smooth tame Frechet manifold.
3. Linear Analysis I
In this section, we prove the existence of $C^{\infty}$ solutions to the linearized problem (2.10) for arbitrarily prescribed smooth target data in $T({\mathcal{T}}^{H})$, in the localized setting of §2.2. The global version of the results here are proved in §5.
Throughout this section, the local rescaled metric $\widetilde{g}$ will be denoted by $g$. Also, since most of the analysis concerns the boundary behavior, the unit normal $\nu_{{\mathcal{C}}}$ to ${\mathcal{C}}$ is generally denoted simply by $\nu$. It is extended into bulk region $\widetilde{U}$ as
$\nu_{{\mathcal{C}}}=\frac{1}{|\nabla x^{1}|}\nabla x^{1}=\frac{1}{|\nabla x^{1}|}g^{\alpha 1}\partial_{\alpha}.$
Clearly, $\nu=\nu_{{\mathcal{C}}}$ is close to its Minkowski corner value, given by (2.7).
3.1. Boundary data equations.
The boundary data in the target space ${\mathcal{T}}^{H}$ consists of the data $([g_{{\mathcal{C}}}],\mu_{g},V_{{\mathcal{C}}})$. The linearizations of these terms are given by
(3.1) $[g_{{\mathcal{C}}}]^{\prime}_{h}=\ring{h}=h^{\mathsf{T}}-\frac{\mathrm{tr}_{{\mathcal{C}}}h^{\mathsf{T}}}{n}g_{{\mathcal{C}}},$
(3.2) $(\mu_{g})^{\prime}_{h}=\tfrac{1}{2}(h_{\nu\nu}+\tfrac{n-2}{n}\mathrm{tr}_{{\mathcal{C}}}h^{\mathsf{T}})\mu_{g},$
(3.3) $(V^{\prime}_{h})^{\mathsf{T}}=-\nabla_{\nu}h(\nu)^{\mathsf{T}}+\delta_{{\mathcal{C}}}\ring{h}+\tfrac{1}{2}d(h_{\nu\nu}+\tfrac{n-2}{n}u)-(A+H\gamma)h(\nu)^{\mathsf{T}}-(\langle D^{2}x^{\alpha},h\rangle\partial_{\alpha})^{\mathsf{T}},$
(3.4) $(V^{\prime}_{h})(\nu)=-\tfrac{1}{2}\nu(h_{\nu\nu})+\tfrac{1}{2}\nu(\mathrm{tr}_{{\mathcal{C}}}h^{\mathsf{T}})+\delta_{{\mathcal{C}}}(h(\nu)^{\mathsf{T}})-h_{\nu\nu}H+\langle A,h^{\mathsf{T}}\rangle-\langle D^{2}x^{\alpha},h\rangle g_{\nu\alpha}.$
Here the superscript $\mathsf{T}$ denotes the component tangent to the boundary ${\mathcal{C}}$. The derivations of (3.1) and (3.2) are elementary while the equations (3.3)-(3.4) follow from the linearization of $V$ in (2.19). Namely, as noted in [2], straightforward computation from (2.19) gives
$(V^{\prime}_{h})^{\mathsf{T}}=-\nabla_{\nu}h(\nu)^{\mathsf{T}}+\beta_{{\mathcal{C}}}h{{}^{\mathsf{T}}}+\tfrac{1}{2}dh_{\nu\nu}-(A+H\gamma)h(\nu)^{\mathsf{T}}-(\langle D^{2}x^{\alpha},h\rangle\partial_{\alpha})^{\mathsf{T}},$
Write $\beta_{{\mathcal{C}}}h^{\mathsf{T}}=\delta_{{\mathcal{C}}}(\ring{h}+\frac{\mathrm{tr}_{{\mathcal{C}}}h^{\mathsf{T}}}{n}g_{{\mathcal{C}}})+\frac{1}{2}d\mathrm{tr}_{{\mathcal{C}}}h^{\mathsf{T}}=\delta_{{\mathcal{C}}}\ring{h}+\frac{n-2}{2n}du$. Substituting this in the equation above gives (3.3). The reason for the choice of the boundary term $\mu_{g}$ is exactly the appearance of $(\mu_{g})^{\prime}_{h}$ in (3.3). Similar computation gives (3.4).
For later purposes, we rewrite (3.3) as
(3.5) $\nabla_{\nu}h(\nu)^{\mathsf{T}}=\widetilde{Y}({\mathcal{T}})+\widetilde{E}(h),$
where $\widetilde{Y}({\mathcal{T}})=-(V^{\prime}_{F})^{\mathsf{T}}+\delta\ring{h}+\frac{1}{2}(h_{\nu\nu}+\frac{n-2}{n}\mathrm{tr}_{{\mathcal{C}}}h^{\mathsf{T}})$ is fully determined by the target data and $\widetilde{E}(h)=-(A+H\gamma)h(\nu)^{\mathsf{T}}-(\langle D^{2}x^{\alpha},h\rangle\partial_{\alpha})^{\mathsf{T}}$ is regarded later as a small error term.
The target data $(\ring{h},\mu^{\prime}_{h})\in T({\mathcal{T}}^{H})$ represent Dirichlet boundary data for $h$ on ${\mathcal{C}}$, while the terms $V^{\prime}_{h}$ represent (to leading order) Neumann boundary data, together with given target data. Note that in these equations, there is no simple equation for the variation of the conformal factor
$u=\mathrm{tr}_{{\mathcal{C}}}h^{\mathsf{T}}.$
An explicit equation for $u$ was derived in [2] via the linearization of the Hamiltonian constraint (2.11). This is given by the following Lemma; we refer to [2] for the detailed proof.
Lemma 3.1.
The function $u$ satisfies the wave equation
(3.6) $-\tfrac{n-1}{n}\Box_{\mathcal{C}}u-\tfrac{1}{n}R_{\mathcal{C}}u=Y({\mathcal{T}})+E(h),$
along the boundary ${\mathcal{C}}$, where
(3.7) $\begin{split}Y({\mathcal{T}})&=-R^{\prime}_{\ring{h}}+\mathrm{tr}_{g}(F-\delta^{*}V^{\prime}_{F})-2(F-\delta^{*}V^{\prime}_{F})(\nu,\nu),\\ E(h)&=-2\langle A^{\prime}_{h},A\rangle+2H_{{\mathcal{C}}}H^{\prime}_{h}+2\langle A\circ A,h\rangle-\mathrm{tr}_{g}[\delta^{*}(W^{\prime}_{h})+(\delta^{*})^{\prime}_{h}V]+2[\delta^{*}(W^{\prime}_{h})+(\delta^{*})^{\prime}_{h}V](\nu,\nu)\\ &\quad-4\mathrm{Ric}(\nu,\nu^{\prime}_{h})-\langle\mathrm{Ric},h\rangle.\end{split}$
The initial data $u$ and $\partial_{t}u$ for $u$ at the initial surface $\Sigma$ are determined by initial and boundary data in $T({\mathcal{T}}^{H})$.
3.2. Construction of local linearized solutions.
In this subsection, we construct solutions $h$ to the linear problem $D\Phi_{g}^{H}(h)=\tau^{\prime}$ for arbitrary $\tau^{\prime}\in T({\mathcal{T}}^{H})$. This is done by an iteration process similar to that used in [2, 3]. Thus, let
(3.8) $\tau^{\prime}=(F,(\gamma_{S}^{\prime},\kappa^{\prime},\nu^{\prime},V_{S}^{\prime}),(\ring{\gamma^{\prime}},\mu^{\prime},V_{{\mathcal{C}}}^{\prime}))$
be a general element in $T_{\tau}({\mathcal{T}}^{H})$, $\tau=\Phi^{H}(g)$. The main result for this section is the following surjectivity or existence theorem.
Theorem 3.2.
There exists $\varepsilon_{0}>0$ such that, for any smooth metric $g$ on $\widetilde{U}$ which is $C^{\infty}$ $\varepsilon$-close, $\varepsilon\leq\varepsilon_{0}$, to a standard Minkowski corner metric $g_{\alpha_{0}}$ as in (2.5), the linearization $D\Phi_{g}^{H}$ at $g$ satisfies: given any $C^{\infty}$ target data $\tau^{\prime}\in T_{\tau}({\mathcal{T}}^{H})$ on $\widetilde{U}$, there exists a variation $h\in T_{g}Met(\widetilde{U})$, such that
(3.9) $D\Phi_{g}^{H}(h)=\tau^{\prime}.$
Thus, the equation
(3.10) $L(h)=F,$
has a $C^{\infty}$ solution $h$ such that
(3.11) $(g_{S},K)^{\prime}_{h}=(\gamma_{S}^{\prime},\kappa^{\prime}),\ \ (\nu_{S})^{\prime}_{h}=\nu^{\prime},\ \ V^{\prime}_{h}=V_{S}^{\prime}\ \mbox{ on }S,$
and
(3.12) $(([g_{\mathcal{C}}])^{\prime}_{h},\mu^{\prime}_{h})=(\ring{\gamma^{\prime}},\mu^{\prime}),\ \ V^{\prime}_{h}=V_{\mathcal{C}}^{\prime}\ \mbox{ on }{\mathcal{C}}.$
Further the constructed solution $h$ depends smoothly on the data $\tau^{\prime}$.
Proof.
Throughout the following $h$ is a symmetric bilinear form, written in adapted local coordinates $x^{\alpha}=(t,x^{i})$ as
$h=h_{\alpha\beta}dx^{\alpha}dx^{\beta}.$
The full variation $h$ is decomposed into a collection of terms as follows. On the boundary ${\mathcal{C}}$, $h^{\mathsf{T}}$ has the form
$h^{\mathsf{T}}=h_{ij}dx^{i}dx^{j},$
where $i,j=0,2,\dots,n$. Then $h^{\mathsf{T}}$ is extended into the bulk $\widetilde{U}$ to have the same form, i.e. without any $dx^{1}\cdot dx^{\alpha}$ term. Similarly, $h(\nu,\cdot)$ is defined on $\widetilde{U}$ by
(3.13) $h(\nu,\cdot)=\frac{1}{|\nabla x^{1}|}g^{1\alpha}h(\partial_{\alpha},\cdot),$
which is defined by the background metric $g$ on $\widetilde{U}$.
In the following, we will work with separate equations for each of the terms
(3.14) $h^{\mathsf{T}},\ h_{\nu\nu},\ h(\nu)^{\mathsf{T}},$
comprising $h$.
0. In this step, we solve the relevant equations with the coupling terms set to zero, i.e. the $P$-terms in the bulk and the $E$-terms in (3.6) and (3.5) on ${\mathcal{C}}$ are set to zero. Thus, the solution is constructed essentially depending only on the given target data in $T({\mathcal{T}}^{H})$. This is done step-by-step for each of the terms in (3.14),
To begin, solve the evolution equation (3.6) for $u$ along ${\mathcal{C}}$, with $Y$ given by the target data $\tau^{\prime}$ and $E=0$ and with given initial conditions at $\Sigma$. As discussed in §2.4, the initial conditions $u|_{\Sigma}$ and $\partial_{t}u|_{\Sigma}$ are determined by the target data in $T({\mathcal{T}})$. Solving (3.6) on ${\mathcal{C}}$ with this data gives the ($C^{\infty}$) Dirichlet boundary value for $u$ along ${\mathcal{C}}$. Given this Dirichlet boundary data, solve next the bulk tangential equation
$-\tfrac{1}{2}\Box_{g}h^{\mathsf{T}}=F^{\mathsf{T}},$
with the given initial conditions on $\widetilde{U}$, as discussed in §2.4 and Lemma 2.1.
Turning next to $h_{\nu\nu}$, the target data $\mu^{\prime}=\frac{1}{2}(h_{\nu\nu}+\frac{n-2}{n}u)\mu_{g}$ and $u$ above determine the Dirichlet boundary value of $h_{\nu\nu}$ on ${\mathcal{C}}$ and so we solve the bulk equation
$-\tfrac{1}{2}\Box_{g}h_{\nu\nu}=F_{\nu\nu},$
with these given Dirichlet boundary conditions for $h_{\nu\nu}$ and with the given initial conditions from $\tau^{\prime}$.
Finally, referring to (3.5), with $\widetilde{Y}$ determined by the target data $\tau^{\prime}$ and setting $\widetilde{E}=0$, the equation (3.5) gives Neumann boundary data $\nabla_{\nu}h(\nu)^{\mathsf{T}}$ for $h(\nu)^{\mathsf{T}}$ and so we solve
$-\tfrac{1}{2}\Box_{g}h(\nu)^{\mathsf{T}}=F(\nu)^{\mathsf{T}},$
in $\widetilde{U}$ with this given Neumann boundary data and with initial data given by the target data in $T({\mathcal{T}})$,
Summing up the solutions above gives a solution $h_{0}$ to
$-\tfrac{1}{2}\Box_{g}h_{0}=F,$
in $\widetilde{U}$ with the prescribed initial and boundary conditions from $\tau^{\prime}$.
Next, define the gauge field $V^{\prime}_{0}$ by $V^{\prime}_{0}=V^{\prime}_{F}+W^{\prime}_{h_{0}}$ where $V^{\prime}_{F}$ and $W^{\prime}_{h_{0}}$ are given as in (2.29) and (2.30) respectively.
The next stages are error, so $O(\varepsilon)$ adjustments to $h_{0}$, so call these now $E$ terms, i.e. $E_{1}$, $E_{2}$, etc.
1. First, solve the evolution equation (3.6) for $u$ along ${\mathcal{C}}$ with $Y=0$, $E=E(h_{0})$ and zero initial data; for the gauge term in (3.7) we use the data $W^{\prime}_{h_{0}}$. This gives Dirichlet boundary data for $u:=u_{1}$ on ${\mathcal{C}}$. Since $E(h_{0})=O(\varepsilon)$ in (3.7), $u_{1}=O(\varepsilon)$ in $C^{\infty}$ on ${\mathcal{C}}$. Using this, we then solve the bulk equation
$-\tfrac{1}{2}\Box_{g}E_{1}^{\mathsf{T}}=-P^{\mathsf{T}}(h_{0}),$
in $\widetilde{U}$ with zero initial data and Dirichlet boundary data given by $h^{\mathsf{T}}=\frac{u_{1}}{n}g_{{\mathcal{C}}}$ on ${\mathcal{C}}$. By construction, $E_{1}^{\mathsf{T}}=O(\varepsilon)$ again by (2.5).
Next, for $(E_{1})_{\nu\nu}$, setting $\mu_{1}=0$, we use the data $(E_{1})_{\nu\nu}=-\frac{n-2}{n}u_{1}$ as Dirichlet boundary data and solve the IBVP
$-\tfrac{1}{2}\Box_{g}(E_{1})_{\nu\nu}=-P_{\nu\nu}(h_{0}),$
in $\widetilde{U}$, with zero initial data and with the Dirichlet data given above.
Finally, solve the IBVP
$-\tfrac{1}{2}\Box_{g}E_{1}(\nu)^{\mathsf{T}}=-P(\nu)^{\mathsf{T}}(h_{0})$
in $\widetilde{U}$, with zero initial data and with Neumann boundary data given by (3.5) with $\widetilde{Y}=0$ and $\widetilde{E}=\widetilde{E}(h_{0})$.
Summing these components together then gives $E_{1}$ solving
$-\tfrac{1}{2}\Box_{g}E_{1}=-P(h_{0}),$
in $\widetilde{U}$. The term $E_{1}$ has zero target initial data and zero target boundary data. As noted above, the boundary data for $E_{1}$ is $O(\varepsilon)$ in $C^{\infty}$, so that
$E_{1}=O(\varepsilon)$
in $C^{\infty}$.
Let $h_{1}=h_{1}+E_{1}$. For the gauge term, we set $V_{1}^{\prime}=V_{0}^{\prime}+W^{\prime}_{E_{1}}$.
2. Next, let $h_{2}=h_{1}+E_{2}$, where $E_{2}$ is constructed in the same way as $E_{1}$ above, with $E_{1}$ in place of $h_{0}$ and $E_{2}$ in place of $E_{1}$. Thus, $E_{2}$ solves
$-\tfrac{1}{2}\Box_{g}E_{2}=-P(E_{1}),$
in $\widetilde{U}$, with zero target initial data and zero target boundary data. The boundary data for $E_{2}$ is constructed by setting $E=E(E_{1})$ in (3.6) and $\widetilde{E}=\widetilde{E}(E_{1})$ in (3.5). It follows as above that
$E_{2}=O(\varepsilon^{2}).$
Set $V_{2}^{\prime}=V_{1}^{\prime}+W^{\prime}_{E_{2}}$.
Continuing inductively in this way, gives the sequence $\{E_{m}\}$, $m=1,2,\cdots$, satisfying
$E_{m}=O(\varepsilon^{m}),$
in $C^{\infty}$ and hence the sequence $h_{k}=h_{0}+\sum_{1}^{k}E_{i}$ converges to a limit $\bar{h}\in C^{\infty}$. By construction
$-\tfrac{1}{2}\Box_{g}(h_{0}+\sum E_{i})=F-P(h_{0}+\sum E_{i}),$
so that
(3.15) $L(\bar{h})=F.$
Hence $\bar{h}$ solves (3.10). The sequence of gauge fields $V_{i}^{\prime}$ also converges in $C^{\infty}$ to a limit field $\bar{V}^{\prime}$. By construction, the pair $(\bar{h},\bar{V}^{\prime})$ satisfies the initial conditions and also satisfies the first boundary condition in (3.12). It remains to show that
$V^{\prime}_{\bar{h}}=\bar{V}^{\prime},$
since $\bar{V}^{\prime}|_{{\mathcal{C}}}=V^{\prime}_{{\mathcal{C}}}$ by construction.
Clearly $V^{\prime}_{\bar{h}}$ and $\bar{V}^{\prime}$ both satisfy (2.27) and have the same initial conditions; the issue is then the Dirichlet boundary data. Now the construction of the Dirichlet boundary value of $\bar{h}$ directly uses the tangential component $(V^{\prime}_{h})^{{\mathcal{T}}}$ of the gauge field $V^{\prime}_{h}$ at ${\mathcal{C}}$; thus the equation (3.3) is directly used as boundary data for the iterative construction of $\bar{h}$. It is only the equation (3.4) for the normal component $V^{\prime}_{h}(\nu)$ that does not enter the construction of $\bar{h}$.
To address this, from the iterative construction of the component $u$ of $\bar{h}$, it follows by taking limit of equation (3.6) that
(3.16) $\begin{split}-\tfrac{n-1}{n}\Box_{\mathcal{C}}\bar{u}-\tfrac{1}{n}R_{\mathcal{C}}\bar{u}=&R^{\prime}_{\ring{h}}-\mathrm{tr}(F-\delta^{*}\bar{V}^{\prime})+2(F-\delta^{*}\bar{V}^{\prime})(\nu,\nu)\\ &+2\langle A^{\prime}_{\bar{h}},A\rangle-2H_{\mathcal{C}}H^{\prime}_{\bar{h}}-2\langle\widehat{A}\circ A,\bar{h}\rangle+4\mathrm{Ric}(\nu,\nu^{\prime}_{\bar{h}})+\langle\mathrm{Ric},\bar{h}\rangle\\ &+\mathrm{tr}[(\delta^{*})^{\prime}_{\bar{h}}V]-2[(\delta^{*})^{\prime}_{\bar{h}}V](\nu,\nu),\\ \end{split}$
along ${\mathcal{C}}$. On the other hand, by (3.15), $(\mathrm{Ric}+\delta^{*}V)^{\prime}_{\bar{h}}=F$ and applying the linearization of the Hamiltonian constraint shows that (3.16) holds with $V^{\prime}_{\bar{h}}$ in place of $\bar{V}^{\prime}$. It follows that
$\mathrm{tr}(\delta^{*}\bar{V}^{\prime})-2\delta^{*}\bar{V}^{\prime}(\nu,\nu)=\mathrm{tr}(\delta^{*}V^{\prime}_{\bar{h}})-2\delta^{*}V^{\prime}_{\bar{h}}(\nu,\nu)\mbox{ on }{\mathcal{C}}.$
Since, as noted above, $(\bar{V}^{\prime})^{\mathsf{T}}=(V^{\prime}_{\bar{h}})^{\mathsf{T}}$, this gives
$\nu\langle\bar{V}^{\prime},\nu\rangle=\nu\langle V^{\prime}_{\bar{h}},\nu\rangle.$
Since both $\bar{V}^{\prime}$ and $V^{\prime}_{\bar{h}}$ satisfy the same gauge equation (2.27), it follows from uniqueness of solutions to the Neumann boundary value problem that $\bar{V}^{\prime}=V^{\prime}_{\bar{h}}$.
Also, the construction is clearly smooth in the data $T(\widetilde{\mathcal{T}}^{H})$ and in $g$. Setting $\bar{h}=h$ completes the proof.
The iteration process above loses derivatives at each stage, due to the loss of derivative in solving the Neumann problem in (3.3) to obtain the solutions $h_{0}(\nu)^{\mathsf{T}},E_{1}(\nu)^{\mathsf{T}},E_{2}(\nu)^{\mathsf{T}}$, etc.
4. Energy estimates
In this section, we derive (apriori) energy estimates for smooth solutions of the linear system $L(h)=F$ with linearized boundary data $(\ring{h},\mu_{h}^{\prime})$, in the local setting of §2.2. Using this, energy estimates in the global setting are then given in §5.
Define a Sobolev $H^{s}$ norm on the target space data in $T_{\tau}({\mathcal{T}}^{H})$ as follows: for $\tau^{\prime}$ as in (3.8),
(4.1) $\begin{split}||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}=[||F||_{H^{s}(\widetilde{U})}&+[||\ring{\gamma^{\prime}}||_{H^{s+1}({\mathcal{C}})}+||\mu^{\prime}||_{H^{s+1}({\mathcal{C}})}+||V^{\prime}_{{\mathcal{C}}}||_{H^{s}({\mathcal{C}})}]\\ &+[||\gamma^{\prime}_{S}||_{H^{s+1}(S_{0})}+||\kappa^{\prime}||_{H^{s}(S_{0})}+||\nu^{\prime}||_{H^{s}(S_{0})}+||V^{\prime}_{S}||_{H^{s}(S_{0})}].\end{split}$
On the right in (4.1), the top line consists of bulk and boundary terms, while the bottom line consists of initial data terms. Note the shift in the derivative index $s$ on the target compared with the discussion in (2.2) or (2.10). Of course $S_{0}$ and ${\mathcal{C}}$ are the corresponding local domains in $\widetilde{U}$.
The main result of this section is the following.
Theorem 4.1.
For the linearization $D\Phi_{g}^{H}$ of $\Phi^{H}$ in (2.10) with $C^{\infty}$ data localized on $\widetilde{U}\supset U$ as in §2.2, one has the tame energy estimate
(4.2) $||h||_{{\mathcal{N}}^{s}(S_{t})}\leq C||D\Phi_{g}(h)||_{H^{s}({\mathcal{T}}^{H})},$
where $C$ is a constant depending only on $g$. In particular,
$KerD\Phi_{g}^{H}=0.$
Proof.
First decompose $h$ into its components as in (3.14). The bulk equation (2.21) then becomes the following collection of equations:
(4.3) $-\tfrac{1}{2}\Box_{g}h^{\mathsf{T}}+P^{\mathsf{T}}(h)=F^{\mathsf{T}},$
(4.4) $-\tfrac{1}{2}\Box_{g}h_{\nu\nu}+P_{\nu\nu}(h)=F_{\nu\nu},$
(4.5) $-\tfrac{1}{2}\Box_{g}h(\nu)^{\mathsf{T}}+P(\nu)^{\mathsf{T}}(h)=F(\nu)^{\mathsf{T}},$
By Lemma 2.1, solutions of these equations satisfy the boundary stable energy estimate (2.23) for Dirichlet boundary data. Here the coupling term $P(h)$ is moved to the right side of each equation and treated as an inhomogeneous bulk term. We assume that $g$ and all target data are $C^{\infty}$ and the target data have compact support in $\widetilde{U}$, away from ${\mathcal{C}}\cap U$, cf. §2.2. Let $\tau^{\prime}=D\Phi^{H}_{g}(h)$.
Recall that $h^{\mathsf{T}}=\ring{h}+\frac{u}{n}g_{{\mathcal{C}}}$ and that the Dirichlet boundary data for $\ring{h}$ is given target data. Next the evolution or wave equation (3.6) for $u$ along ${\mathcal{C}}$ gives the following energy estimate for $u$ along ${\mathcal{C}}$:
(4.6) $\begin{split}||u||_{\widetilde{H}^{s}(\Sigma_{t})}\leq C&[||F||_{H^{s-1}({\mathcal{C}}_{t})}+||\ring{h}||_{H^{s+1}({{\mathcal{C}}_{t})}}+||V^{\prime}_{h}||_{\bar{H}^{s}({\mathcal{C}}_{t})}+\varepsilon||h||_{\bar{H}^{s}({\mathcal{C}}_{t})}]\\ +&\ C[||u||_{H^{s}(\Sigma_{0})}+||\partial_{t}u||_{H^{s-1}(\Sigma_{0})}].\end{split}$
Here $\widetilde{H}^{s}(\Sigma_{t})$ denotes the Sobolev norm with both $t$-derivatives as well as derivatives tangent to $\Sigma_{t}$, but no derivatives normal to ${\mathcal{C}}$. The extra derivative required on $\ring{h}$ comes from the $(R_{{\mathcal{C}}})^{\prime}_{\ring{h}}$ term in (3.7). Except for terms involving $\delta^{*}V^{\prime}_{h}$, the right-side of (3.6), i.e. the inhomogeneous term, is of the form $\varepsilon(\partial h)$, giving rise to the $\varepsilon$ term in (4.6). Note all derivatives, including normal derivatives along ${\mathcal{C}}$ are included in the $\varepsilon$-term in (4.6).
The last line in (4.6) consists of initial data and is bounded by
$||u||_{H^{s}(\Sigma_{0})}+||\partial_{t}u||_{H^{s-1}(\Sigma_{0})}\leq C[||\gamma^{\prime}_{S}||_{H^{s+1/2}(S_{0})}+||\kappa^{\prime}||_{H^{s-1/2}(S_{0})}],$
where we have used the Sobolev trace theorem and Remark 2.4 and Remark 2.7. Similarly,
$||F||_{H^{s-1}({\mathcal{C}})}\leq C||F||_{H^{s-1/2}(\widetilde{U})}.$
Next from the gauge wave equation (2.27) and the boundary stable energy estimate with Dirichlet boundary data (2.23) we obtain
(4.7) $\begin{split}||V^{\prime}_{h}||_{\bar{H}^{s}({\mathcal{C}}_{t})}\leq C&[||F||_{H^{s}(\widetilde{U})}+||V^{\prime}_{{\mathcal{C}}}||_{H^{s}({\mathcal{C}}_{t})}+\varepsilon||h||_{H^{s}(\widetilde{U})}]\\ +&C[||V^{\prime}_{h}||_{H^{s}(S_{0})}+||\partial_{t}V^{\prime}_{h}||_{H^{s-1}(S_{0})}].\end{split}$
Note that this estimate requires an extra derivative on the bulk $F$ term, due to the $\beta(F)$ term in (2.27). Also, while $V^{\prime}_{h}|_{S_{0}}=V^{\prime}_{S}$ is given target data, the term $\partial_{t}V^{\prime}_{h}|_{S_{0}}$ is not. To obtain a bound on this term, we use the linearized constraint equations (2.13)-(2.14); this gives
$||\partial_{t}V^{\prime}_{h}||_{H^{s-1}(S_{0})}\leq C[||h_{S}||_{H^{s+1}(S_{0})}+||K^{\prime}_{h}||_{H^{s}(S_{0})}+||F||_{H^{s-1}(S_{0})}+||V^{\prime}_{h}||_{H^{s}(S_{0})}]\leq C||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}.$
Given this control on the Dirichlet boundary value of $h^{\mathsf{T}}$, (4.3) and (2.23) then gives the estimate
(4.8) $\begin{split}||h^{\mathsf{T}}||_{\bar{\mathcal{H}}^{s}(S_{t})}&\leq C[||h^{\mathsf{T}}||_{H^{s}({\mathcal{C}}_{t})}+\varepsilon||h||_{H^{s}(\widetilde{U}_{t})}+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}]\\ &\leq C[\int_{0}^{t}[||\ring{h}||_{H^{s+1}({{\mathcal{C}}_{t^{\prime}})}}+||V^{\prime}_{{\mathcal{C}}}||_{H^{s}({\mathcal{C}}_{t^{\prime}})}.+\varepsilon||h||_{\bar{H}^{s}({\mathcal{C}}_{t^{\prime}})}]dt^{\prime}+\varepsilon||h||_{H^{s}(\widetilde{U}_{t})}+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}],\\ \end{split}$
where the last inequality comes from integration of (4.6)-(4.7) over $t^{\prime}\in[0,t]$. Taking the maximum over $t\in I$ then gives
(4.9) $\max_{t\in[0,1]}||h^{\mathsf{T}}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C[||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}+\varepsilon||h||_{\widehat{H}^{s}(\widetilde{U})}],$
where $||h||_{\widehat{H}^{s}(\widetilde{U})}:=||h||_{H^{s}(\widetilde{U})}+||h||_{\bar{H}^{s}({\mathcal{C}})}$.
Similarly, the Dirichlet energy estimate for $h_{\nu\nu}$ gives
$\max_{t\in[0,1]}||h_{\nu\nu}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C[||h_{\nu\nu}||_{H^{s}({{\mathcal{C}})}}+\varepsilon||h||_{H^{s}(\widetilde{U})}+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}],$
and so using the definition of $\mu^{\prime}_{h}$, $h_{\nu\nu}=\mu^{\prime}_{h}-\frac{n-2}{n}u$, together with (4.9) gives
(4.10) $\max_{t\in[0,1]}||h_{\nu\nu}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C[||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}+\varepsilon||h||_{\widehat{H}^{s}(\widetilde{U})}],$
Finally for $h(\nu)^{\mathsf{T}}$, from the Neumann boundary estimate (2.24), we have
$||h(\nu)^{\mathsf{T}}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C[||\nabla_{\nu}h(\nu)^{\mathsf{T}}||_{H^{s-1/3}({{\mathcal{C}}_{t})}}+\varepsilon||h||_{H^{s}(\widetilde{U}_{t})}+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}].$
Examination of the form of the Neumann boundary data in (3.3) gives
(4.11) $\begin{split}||h(\nu)^{\mathsf{T}}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq&C[||\ring{h}||_{H^{s+2/3}({\mathcal{C}}_{t})}+||\mu^{\prime}_{h}||_{H^{s+2/3}({\mathcal{C}}_{t})}\\ &+||V^{\prime}_{{\mathcal{C}}}||_{H^{s-1/3}({\mathcal{C}})}+\varepsilon(||h||_{H^{s-1/3}({\mathcal{C}}_{t})}+||h||_{H^{s}(\widetilde{U}_{t})})+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}].\end{split}$
Summing up the estimates (4.6)-(4.11), by choosing $\varepsilon$ sufficiently small depending on $g$, one may absorb the $\varepsilon||h||_{\bar{H}^{s}(\widetilde{U})}$ terms on the right side of these equations into the total sum on the left, using the fact that $|h||_{\bar{H}^{s}(\widetilde{U})}\leq\max_{t\in[0,1]}|h||_{\bar{\mathcal{H}}^{s}(S_{t})}$, giving then
(4.12) $\max_{t\in[0,1]}||h||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}.$
We note that the estimate (4.11) is the key ingredient in establishing the bound (4.12) in comparison with the work in [2, 3].
The bound on the left in (4.12) gives a bound on
(4.13) $h\in L^{\infty}(I,\bar{\mathcal{H}}^{s}(S)).$
However, from the regularity theory for wave equations with Dirichlet boundary data as in Lemma 2.1, by performing the computations in the proof of Theorem 4.1 over arbitrarily small time intervals $I^{\prime}=[t_{0},t_{1}]\subset I$, it is easy to verify that this can be improved to a bound on
$h\in C^{0}(I,\bar{\mathcal{H}}^{s}(S)),$
so that (4.12) gives in fact
(4.14) $||h||_{{\mathcal{N}}^{s}}^{2}\leq C||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}.$
This completes the proof of Theorem 4.1.
5. Linear Analysis II
In this section, we prove the global version of the results of §3 and §4. This is done by patching together local solutions and estimates to obtain global in space solutions and estimates. The following result proves the well-posedness of the IBVP for the linearized problem, allowing for loss-of-derivatives.
Theorem 5.1.
Let $(M,g)$ be a smooth globally hyperbolic Lorentzian manifold, $M\simeq I\times S$, with compact Cauchy surface $S$ having compact boundary $\partial S=\Sigma$ and with timelike boundary $({\mathcal{C}},g_{{\mathcal{C}}})$. Then for any smooth $\tau^{\prime}\in T_{\tau}({\mathcal{T}}^{H})$, the equation
(5.1) $D\Phi_{g}^{H}(h)=\tau^{\prime}$
has a unique $C^{\infty}$ solution $h$. The constructed solution $h$ depends smoothly on the data $\tau^{\prime}$.
Moreover, any $h$ satisfying (5.1) satisfies the tame energy estimate
(5.2) $||h||_{{\mathcal{N}}^{s}}\leq C||D\Phi_{g}^{H}(h)||_{H^{s}(T^{H})},$
where $C$ depends only on $g$.
Proof.
Given the previous local results, the proof of Theorem 5.1 is essentially standard. We provide the details below, following the proof of the analogous result in [2].
Given $(M,g)$ as in Theorem 5.1, first choose an open cover $\{U_{i}\}_{i=1}^{N}$ of the corner $\Sigma$, where each $U_{i}$ is sufficiently small so that the local existence result Theorem 3.2 holds for each $U_{i}$. Also, choose an open set $U_{0}$ in the interior $M\setminus{\mathcal{C}}$ such that $(\cup_{i=1}^{N}U_{i})\cup U_{0}$ also covers a tubular neighborhood of the initial surface, i.e. $S\times[0,t_{0}]$ for some time $t_{0}>0$.
Let $\{\widetilde{U}_{i}\}_{i=0}^{N}$ be a thickening of the cover $\{U_{i}\}_{i=0}^{N}$, so that $U_{i}\subset\widetilde{U}_{i}$, as described in §2.2. Let $\rho_{i}$ be a partition of unity subordinate to the cover $\{\widetilde{U}_{i}\}_{i=0}^{N}$, so that ${\rm supp}\rho_{i}\subset\widetilde{U}_{i}$, $\rho_{i}=1$ on $U_{i}$ and $\sum_{i}\rho_{i}=1$. Let
$\tau^{\prime}=\big{(}F,(\gamma^{\prime},\kappa^{\prime},\nu^{\prime},V_{S}^{\prime}),(\ring{\gamma^{\prime}},\eta^{\prime},V_{{\mathcal{C}}}^{\prime})\big{)},$
be arbitrary $C^{\infty}$ data in $T({\mathcal{T}}^{H})$ on $M$. Then the data $\rho_{i}\tau^{\prime}$ has compact support in the sense of §2.2 in $\widetilde{U}_{i}$ and by Theorem 3.2 there exists a solution $h_{i}$ in $\widetilde{U}_{i}$ satisfying
(5.3) $D\Phi^{H}_{g}(h_{i})=\rho_{i}\tau^{\prime}.$
(Here $\tau^{\prime}$ is rescaled by the relevant powers of $\lambda$ as in §2.2, but the notation will not reflect this). Moreover, as noted at the end of §2.2, by the finite propagation speed property, the solution $h_{i}$ has compact support in the sense of §2.2 for $t\in[0,t_{i}]$ for some $t_{i}>0$. Thus, $h_{i}$ extends smoothly as the zero solution on $M_{t_{i}}\setminus\widetilde{U}_{i}$.
Similarly, in the interior region $U_{0}\subset\widetilde{U}_{0}$, let $h_{0}$ be the solution to the Cauchy problem
$L(h_{0})=\rho_{0}F\ \mbox{ in }\widetilde{U}_{0},\ \ (g_{S},K,\nu_{S},V_{g})^{\prime}_{h}=\rho_{0}(\gamma^{\prime},\kappa^{\prime},\nu^{\prime},V_{S}^{\prime})\ \mbox{ on }S_{0}.$
Then again $h_{0}$ has compact support in $\widetilde{U}_{0}$ for some time interval $[0,t_{0}]$ and as above extends to $M_{t_{0}}$. We relabel so that $t_{0}$ is a common time interval for all $h_{i}$, $i\geq 0$.
The sum
(5.4) $h=\sum_{i=0}^{N}h_{i},$
is thus well-defined on $M_{t_{0}}$ and by linearity
(5.5) $D\Phi_{g}^{H}(h)=\tau^{\prime}.$
(For this global solution, the rescaling above is of course removed). By the local energy estimate in Theorem 4.2, the solution $h_{i}$ of (5.3) is unique. It follows that the global solution $h$ is the unique smooth solution (5.5), since if $h^{\prime}$ is any other solution of (5.5), one may apply the local decomposition (5.4) in the same way to $h^{\prime}$. Moreover, the partition of unity property shows that the local energy estimates from (4.2) sum to give the global energy estimate (5.2).
The smoothness statement follows from that in Theorem 3.2.
6. Nash-Moser theory
The analysis in §5 shows that the linearized IBVP with conformal-volume (or twisted Dirichlet) boundary data is well-posed and has tame energy estimates if one allows for a loss of derivatives for the initial and boundary data. This is the setting of the Nash-Moser implicit function theorem, which we now apply to prove Theorem 1.1.
First we prove a general off-shell result.
Theorem 6.1.
The mapping
(6.1) $\Phi^{H}:Met(M)\rightarrow{\mathcal{T}}^{H}$
$\Phi^{H}(g)=\big{(}\mathrm{Ric}_{g}+\delta^{*}V,(\gamma_{S},K,\nu_{S},V|_{S}),([g_{{\mathcal{C}}}],\mu_{g},V|_{{\mathcal{C}}})\big{)},$
is a smooth tame diffeomorphism between tame Frechet manifolds.
Proof.
By Proposition 2.8, the target space ${\mathcal{T}}^{H}$ is a smooth tame Frechet manifold, as is the domain $Met(M)$. By construction, it is easy to see that the map $\Phi^{H}$ is then a smooth, tame map between Frechet manifolds, cf. [7] for the basic definitions regarding tame maps and Frechet manifolds here.
By Theorem 5.1, the derivative $D\Phi_{g}^{H}$ is a tame isomorphism at any $g\in Met(M)$ and by (5.2), the inverse $(D\Phi_{g}^{H})^{-1}$ exists and is a tame isomorphism. The Nash-Moser inverse function theorem, cf. [7], [14], then implies that $\Phi^{H}$ is everywhere a smooth local diffeomorphism.
Next we claim that $\Phi^{H}$ in (6.1) is not only locally one-to-one but in fact globally one-to-one. To see this, suppose $g_{1},g_{2}$ are two metrics with the same gauged target data, i.e. $\Phi^{H}(g_{1})=\Phi^{H}(g_{2})$. If $g_{1}\neq g_{2}$, then there exists an arbitrarily small neighborhood $U$ of some $p\in\Sigma$ such that $g_{1}\neq g_{2}$ in $U$. One may then perform a localization or rescaling of each $g_{i}$ as in §2.2 so that the almost constant coefficient metrics $\widetilde{g}_{i}$ satisfy $\widetilde{g}_{1}\neq\widetilde{g}_{2}$ in $U$. However, by (2.5), the metrics $\widetilde{g}_{i}$ satisfy
$||\widetilde{g}_{2}-\widetilde{g}_{1}||_{C^{\infty}}\leq\varepsilon,$
in $U$, for a prescribed $\varepsilon>0$. It then follows from the local uniqueness in the Nash-Moser theorem above that
$\widetilde{g}_{2}=\widetilde{g}_{1}$
in some $U^{\prime}\subset U$. Hence $g_{2}=g_{1}$ in $U^{\prime}$, which proves the global uniqueness.
To see that $\Phi^{H}$ is globally surjective, let $\tau\in{\mathcal{T}}^{H}$ be arbitrary. Consider a small corner neighborhood $U$ of $p\in\Sigma$ and rescale the data $\tau$ to $\widetilde{\tau}$. Then $Q\sim 0$, $\gamma_{S}$ is close to a flat half-space $\{x^{1}\leq 0\}$ in $\mathbb{R}^{n}$, $\kappa\sim 0$, $\nu_{S}$ is close to the value in (2.7) and $V\sim 0$. Thus $\widetilde{\tau}$ is close to the value $\widetilde{\tau}_{0}=\Phi(\widetilde{g}_{\alpha_{0}})$, where $\widetilde{g}_{\alpha_{0}}$ is the Minkowski corner metric in (2.6). The Nash-Moser theorem above applies and gives the existence of $\widetilde{g}$ in $U$ such that $\Phi(\widetilde{g})=\widetilde{\tau}$. Rescaling back then gives the existence of $g$ such that $\Phi(g)=\tau$ in $U$. Since $p\in\Sigma$ is arbitrary, such local solutions exist everywhere near $\Sigma$ and the uniqueness above implies the local solutions agree on overlaps. Together with existence and uniqueness of solutions to the Cauchy problem on $S$, this gives the existence of $g$ on $M$ (for small time) with $\Phi(g)=\tau$.
This completes the proof of Theorem 6.1.
Next we turn to the proof of Theorem 1.1
Proof.
Let
${\mathcal{T}}^{H}_{0}=\{\big{(}0,(\gamma,\kappa,\nu_{S},0),(\gamma_{{\mathcal{C}}},\mu,0)\big{)}\}\subset{\mathcal{T}}^{H},$
so $Q$ and the initial and boundary data of the gauge field in ${\mathcal{T}}^{H}$ are set to zero. It is proved in [3] that ${\mathcal{T}}^{H}_{0}$ is a smooth submanifold of ${\mathcal{T}}^{H}$. (The fact that the boundary data here is twisted Dirichlet boundary data as opposed to Dirichlet boundary data plays no role in the proof).
Observe that
$(\Phi^{H})^{-1}({\mathcal{T}}^{H}_{0})=\mathbb{E}^{H}\simeq{\mathcal{E}}_{*}.$
Namely, $\Phi^{-1}({\mathcal{T}}^{H}_{0})$ is the space of Lorentz metrics $g$ on $M$ with $Q=0$ on $M$ and $V=0$ on $S\cup{\mathcal{C}}$. Since then $\mathrm{Ric}_{g}+\delta^{*}V=0$, it follows from Remark 2.4 that $V=0$ on $M$. Thus the metric $g$ is in harmonic gauge and $\mathrm{Ric}_{g}=0$, with given unit timelike normal $\nu_{S}$ at $S$. It follows then from Theorem 6.1 that the map
$\Phi^{H}:\mathbb{E}^{H}\rightarrow{\mathcal{T}}^{H}_{0},$
is a smooth, tame diffeomorphism between Frechet manifolds, with smooth, tame inverse $(\Phi^{H})^{-1}$. This completes the proof of Theorem 1.1.
Next we give the proof of Corollary 1.2.
Proof.
We have shown in the proof of Theorem 1.1 above that $\mathbb{E}^{H}$ is a smooth, tame Frechet manifold. The group $\mathrm{Diff}_{0}(M)/\mathrm{Diff}_{1}(M)$ is a smooth tame Frechet Lie group (cf. [7, I.4.6]) which acts smoothly and freely on $\mathbb{E}^{H}$. It then follows from standard theory, cf. [7, Cor.III.2.5.2] that the quotient $\mathbb{E}^{H}/(\mathrm{Diff}_{0}(M)/\mathrm{Diff}_{1}(M))$ is a smooth, tame Frechet manifold. As noted in §1,
${\mathcal{E}}\simeq\mathbb{E}^{H}/(\mathrm{Diff}_{0}(M)/\mathrm{Diff}_{1}(M)),$
and hence ${\mathcal{E}}$ is a smooth, tame Frechet manifold.
Remark 6.2.
All of the discussion above has been phrased in the $C^{\infty}$ context. However, all the results may be extended to Sobolev spaces $H^{s}$ of finite differentiability. Thus, as discussed in §2, the maps $\Phi$ and $\Phi^{H}$ are well-defined smooth maps between a scale of Banach manifolds,
$\Phi^{H}:Met^{s}(M)\rightarrow({\mathcal{T}}^{H})^{s},$
indexed by $s$. The energy estimates in (5.2) hold on the $H^{s}$ completions of the domain and target spaces for suitable values of $s$, giving the existence of a tame inverse for the derivative $D\Phi^{H}$. One may then apply the finite derivative version of the Nash-Moser inverse function theorem due to Zehnder [13], or [14, Thm.6.3.3]; in the notation used there, the loss of derivative is $\gamma=3$. However, we will not carry out here the details that the hypotheses of [14, Thm.6.3.3] do hold.
Remark 6.3.
As in [2, 3], the results in this work hold for any value of the cosmological constant $\Lambda\in\mathbb{R}$. The results also hold for the Einstein equations coupled to Kaluza-Klein matter fields, which derive from the vacuum Einstein equations in higher dimensions by symmetry reduction. The boundary data of the KK fields are then induced from the higher dimensional vacuum boundary data in ${\mathcal{B}}$.

### v2

@@
 (1.3) $([g_{{\mathcal{C}}}],\mu_{g}),$
-consisting of the conformal class $[g_{{\mathcal{C}}}]$ of the induced boundary metric, together with the scalar
+consisting of the conformal class $g_{{\mathcal{C}}}]$ of the induced boundary metric, together with the scalar
 (1.4) $\mu_{g}=dv_{g}(dv_{g_{{\mathcal{C}}}})^{-2/n}.$
@@
 The product $\mu_{g}$ is viewed as a scalar density with respect to an arbitrarily chosen background volume form on $M$.
-The particular form of $\mu_{g}$ in (1.4) is chosen since it interacts well with the choice of gauge used here - the harmonic or wave coordinate gauge.
+The particular form of $\mu_{g}$ in (1.4) is chosen since it is well adapted to the choice of gauge used here - the harmonic or wave coordinate gauge.
 Different choices of gauge may lead to other expressions for the scalar $\mu_{g}$, but this will not be pursued further here.
-Note that the pair of boundary data $([g_{{\mathcal{C}}}],dv_{g_{{\mathcal{C}}}})$, when taken together, is exactly Dirichlet boundary data $g_{{\mathcal{C}}}$ studied in [3] - which was proved to be ill-posed (i.e.
-not well-posed) in general, i.e.
-without further conditions.
-Thus the boundary data (1.3) can be viewed as a modification of Dirichlet boundary data, involving a twisting by the volume form of $g$.
-The boundary data (1.3) involves no derivatives of the metric at ${\mathcal{C}}$.
+Note that the pair of boundary data $([g_{{\mathcal{C}}}],dv_{g_{{\mathcal{C}}}})$, when taken together, is exactly Dirichlet boundary data $g_{{\mathcal{C}}}$ studied in [3].
+This boundary data was proved to be ill-posed (i.e.
+not well-posed) without further conditions; this is discussed further below.
+The boundary data (1.3) can be viewed as a modification of Dirichlet boundary data, involving a twisting by the volume form of $g$.
+Observe that the boundary data (1.3) involves no derivatives of the metric at $\mathcal{C}$}.
 For $\mathcal{B}$} as in (1.2), there is a natural map to initial and boundary data given by
-(1.5) $\begin{split}&\Phi:\mathbb{E}\rightarrow\mathcal{I}_{0}\times_{c}{\mathcal{B}},\\ &\Phi(g)=(g_{S},K,[g_{{\mathcal{C}}}],\mu_{g}),\\ \end{split}$
+[…]
 where $g_{S}$ is the metric on $S$ induced by $g$ and $K$ is the induced second fundamental form (extrinsic curvature) on $S$.
-The subscript $c$ denotes the compatibility conditions between the initial and boundary data at the corner $\Sigma$.
-Due to the diffeomorphism invariance of the vacuum Einstein equations, it is well-known that the map $\Phi$ cannot be one-to-one and so $\Phi$ cannot be an effective parametrization of $\mathbb{E}$.
+The subscript $c$ denotes the $C^{\infty}$ compatibility conditions between the initial and boundary data at the corner $\Sigma$.
+Due to the diffeomorphism invariance of the vacuum Einstein equations, it is well-known that the map $\Phi$ cannot be one-to-one and so $\Phi$ cannot be an effective parametrization of the space of vacuum solutions $\mathbb{E}$.
 Let $\mathbb{E}^{H}$ be the space of vacuum Einstein metrics in harmonic or wave gauge; cf.
 §2.3 for details.
-Let $\mathrm{Diff}_{1}(M)\subset\mathrm{Diff}_{0}(M)$ be the subgroup of diffeomorphisms $\varphi$ such that $\varphi$ is the identity to first order at $S$; equivalently,
+Let $\mathrm{Diff}_{1}(M)\subset\mathrm{Diff}_{0}(M)$ be the subgroup of diffeomorphisms $\varphi$ such that $\varphi$ is the identity to first order at $S$; equivalently, $\varphi\in\mathrm{Diff}_{0}(M)$ and
 $\varphi_{*}(\partial_{t})=\partial_{t}\ \ {\rm on}\ \ S.$
-There are no $1^{\rm st}$ order restrictions for $\varphi\in\mathrm{Diff}_{0}(M)$ along ${\mathcal{C}}$.
+There are no $1^{\rm st}$ order restrictions for $\varphi\in\mathrm{Diff}_{0}(M)$ along $\mathcal{C}$}, except for the smooth compatibility condition at the corner $\Sigma$.
 For any metric $g$ such diffeomorphisms leave the future time-like normal vector $\nu_{S}$ to $S$ in $(M,g)$ invariant: $\varphi_{*}(\nu_{S})=\nu_{S}$.
@@
 The main result of this paper is the following (local-in-time) well-posedness result for the IBVP with boundary data given by (1.3).
-Let $\mathbb{V}_{S}^{\prime}$ be the space of smooth vector fields along $S$ nowhere tangent to $S$ (the target space for normal vector fields along $S$) and form the product space $\mathcal{I}_{0}\times\mathbb{V}^{\prime}\times{\mathcal{B}}$; let $[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}\subset\mathcal{I}_{0}\times\mathbb{V}^{\prime}\times{\mathcal{B}}$ be the subset consisting of data smoothly compatible at the corner $\Sigma$, cf.
+Let $\mathbb{V}_{S}^{\prime}$ be the space of smooth vector fields along $S$ nowhere tangent to $S$ (the target space for normal vector fields along $S$) and form the product space $\mathcal{I}_{0}\times\mathbb{V}^{\prime}\times{\mathcal{B}}$; let $\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}\subset\mathcal{I}_{0}\times\mathbb{V}^{\prime}\times{\mathcal{B}}$ be the subset consisting of data smoothly compatible at the corner $\Sigma$, cf.
 §2.4 for details.
 Theorem 1.1.
-The spaces $\mathbb{E}^{H}$ and $[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$ are smooth tame Frechet manifolds and the map
-(1.7) $\begin{split}&\Phi^{H}:\mathbb{E}^{H}\rightarrow[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c},\\ &\Phi^{H}(g)=(g_{S},K,\nu_{S},[g_{{\mathcal{C}}}],\mu_{g}),\\ \end{split}$
-is (locally-in-time) a smooth tame diffeomorphism of $\mathbb{E}^{H}$ onto $[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$ with smooth, tame inverse.
+The spaces $\mathbb{E}^{H}$ and $\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$ are smooth tame Frechet manifolds and the map
+[…]
+is (locally-in-time) a smooth tame diffeomorphism of $\mathbb{E}^{H}$ onto $\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$ with smooth, tame inverse.
 Thus the IBVP with boundary data (1.3) is well-posed in harmonic gauge.
-Briefly, vacuum Einstein metrics on $M$, with arbitrarily given compatible initial and boundary data, exist for a short time, and are uniquely determined in harmonic gauge by their initial data in $\mathcal{I}_{0}\times\mathbb{V}_{S}^{\prime}$ and boundary data in ${\mathcal{B}}$.
-Further, such solutions depend continuously and in fact smoothly on the target data in $[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$; the IBVP is stable with respect to variation of the target data.
+Briefly, smooth vacuum Einstein metrics on $M$, with arbitrarily given smoothly compatible initial and boundary data, exist for a short time, and are uniquely determined in harmonic gauge by their initial data in $\mathcal{I}_{0}\times\mathbb{V}_{S}^{\prime}$ and boundary data in $\mathcal{B}$}.
+Further, such solutions depend continuously and in fact smoothly on the target data in $\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$; the IBVP is stable with respect to variation of both the initial data and the boundary data in $\mathcal{B}$}.
 A version of Theorem 1.1 holds also off-shell (for non-vacuum Einstein metrics), cf.
@@
 The marked moduli space $\mathcal{E}$$_{*}$ and the moduli space $\mathcal{E}$} of vacuum Einstein metrics on $M$ as above are smooth tame Frechet manifolds.
-Note that this result does not depend on any choice of boundary data space ${\mathcal{B}}$; it is intrinsic to the structure of the moduli space ${\mathcal{E}}$, independent of any map to boundary data.
+Note that Corollary 1.2 does not depend on any choice of boundary data space $\mathcal{B}$}; it is intrinsic to the structure of the moduli space $\mathcal{E}$}, independent of any map to boundary data.
 The smoothness of the moduli space $\mathcal{E}$} has not been established previously.
-In fact it is known to be false in general for the pure Cauchy problem with compact Cauchy surface $S$ (with $\partial S=\emptyset$); this issue is closely related to the failure of the linearization stability for solutions of the Cauchy problem at Killing initial data, cf.
+In fact it is known to be false in general for the pure Cauchy problem with compact Cauchy surface $S$ with $\partial S=\emptyset$; this issue is closely related to the failure of the linearization stability for solutions of the Cauchy problem at Killing initial data, cf.
 [12].
@@
 In other words, consider the extension of the map $\Phi$ in (1.5) to general (non-gauged) vacuum Einstein metrics:
-(1.8) $\begin{split}&\widehat{\Phi}:\mathbb{E}\rightarrow[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c},\\ &\widehat{\Phi}(g)=(g_{S},K,\nu_{S},[g_{{\mathcal{C}}}],\mu_{g}).\\ \end{split}$
+[…]
 In comparing (1.8) and (1.7), in general,
@@
 The map $\widehat{\Phi}$ is not constant on the fibers of the projection $\pi_{*}$; this is due to the fact that the groups $\mathrm{Diff}_{1}(M)$ and $\mathrm{Diff}_{0}(M)$ act non-trivially on the target data.
-To explain this in more detail, while the boundary data $[g_{{\mathcal{C}}}]$ is invariant under the action of $\mathrm{Diff}_{1}(M)$ (and $\mathrm{Diff}_{0}(M)$), the volume term $\mu_{g}$ transforms as
-(1.10) $(\varphi,\mu_{g})\rightarrow\mu_{\varphi^{*}g}=(\varphi^{*}dv_{g})(dv_{g_{\mathcal{C}}})^{-2/n}.$
+To explain this in more detail, while the boundary data $g_{{\mathcal{C}}}]$ is invariant under the action of $\mathrm{Diff}_{1}(M)$ (and $\mathrm{Diff}_{0}(M)$), the volume term $\mu_{g}$ transforms as
+(1.10) $(\varphi,\mu_{g})\rightarrow\mu_{\varphi^{*}g}=(\varphi^{*}dv_{g})(dv_{g_{\mathcal{C}}})^{-2/n},$
+and generally $\varphi^{*}dv_{g}\neq dv_{g}$ along $\mathcal{C}$}.
 For example, in $4$-dimensions, the boundary degrees of freedom for the metric $g$ require $10$ independent boundary conditions, $4$ of which are associated with the choice of gauge.
-Here, of the remaining $6$ degrees of freedom, $5$ are geometric (corresponding to the specification of $[g_{{\mathcal{C}}}]$), but the scalar volume term $\mu_{g}$ is not (since it is not invariant under the action of $\mathrm{Diff}_{0}(M)$).
-This may be compared with the works in [6], where only $3$ boundary conditions are geometric (in a restricted sense); the geometric nature of the boundary conditions in [8] is less clear.
+Here, of the remaining $6$ degrees of freedom, $5$ are geometric (corresponding to the specification of $g_{{\mathcal{C}}}]$), but the scalar volume term $\mu_{g}$ is not (since it is not invariant under the action of $\mathrm{Diff}_{0}(M)$).
+This may be compared with the works in [7],[9].
+In [7] at most $3$ boundary conditions may be considered geometric in a restricted sense while the geometric nature of the boundary conditions in [9] is less clear.
 Moreover, Theorem 1.1 holds in all dimensions while the works in [7] and [9] are restricted to $4$-dimensions.
-A related but distinct perspective is to return to the map $\Phi^{H}$ in (1.7) and use the identification (1.6) to obtain via Theorem 1.1 the diffeomorphism
-(1.11) $\begin{split}&\Phi_{*}:{\mathcal{E}}_{*}=\mathbb{E}/\mathrm{Diff}_{1}(M)\rightarrow[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c},\\ &\Phi_{*}[g]=(g_{S},K,\nu_{S},[g_{{\mathcal{C}}}],\mu_{g}).\\ \end{split}$
+Theorem 1.1 should be compared with two other recent related results.
+First, in [3], the IBVP was proved to be well-posed for Dirichlet boundary data in an open neighborhood of solutions $(M,g)$ for which the Brown-York stress-energy tensor
+(1.11) $\Pi=Hg_{{\mathcal{C}}}-A,$
+has the same Lorentz signature $(-+\cdots+)$ as the induced metric $g_{{\mathcal{C}}}$ on $\mathcal{C}$}.
+Here $A$ is the second fundamental form of ${\mathcal{C}}\subset(M,g)$ and $H=\mathrm{tr}_{g_{{\mathcal{C}}}}A$ is the mean curvature of the boundary.
+While this is the first result on well-posedness for geometric boundary data, it does not hold in full generality, i.e.
+without some assumption such as (1.11), cf.
+[4], [5] for examples of ill-posedness and further discussion.
+Next, it was recently proved in [10] that the IBVP for (the geometric) conformal mean curvature boundary data $([g_{{\mathcal{C}}}],H)$ is ill-posed at the linearized level at certain backgrounds.
+More precisely, this was proved to be the case when the IBVP is linearized at a standard spherical cylinder in Minkowski $\mathbb{R}^{1,3}$, These remarks illustrate some the difficulties in establishing a general well-posedness result for a geometric IBVP for the Einstein equations, Theorem 1.1 may be viewed as a further step in this direction.
+Returning again to Theorem 1.1, consider the map $\Phi^{H}$ in (1.7) and use the identification (1.6) to obtain via Theorem 1.1 the diffeomorphism
+[…]
 As noted above, the group $\mathrm{Diff}_{0}(M)$ acts on $\mathcal{E}$$_{*}$ with quotient ${\mathcal{E}}={\mathcal{E}}_{*}/\mathrm{Diff}_{0}(M)$ and with fibers given by the orbits of the action of $\mathrm{Diff}_{0}(M)/\mathrm{Diff}_{1}(M)$ on $\mathcal{E}$$_{*}$.
@@
 The map $\Phi_{*}$ is equivariant with respect to these actions and so one obtains an induced map on the quotient spaces:
-(1.12) $\begin{split}&\Phi_{*}:{\mathcal{E}}=\mathbb{E}/\mathrm{Diff}_{0}(M)\rightarrow\mathcal{I}_{0}\times_{c}[{\mathcal{B}}],\\ &\Phi_{*}[g]=(g_{S},K,[g_{{\mathcal{C}}}]).\\ \end{split}$
 […]
+[…]
+Corollary 1.3.
+The map $\Phi_{*}$ in (1.13) is surjective.
+Thus, given arbitrary smooth initial data $(\gamma_{S},\kappa)\in\mathcal{I}_{0}$ together with a smoothly compatible conformal class $\gamma_{{\mathcal{C}}}]\in{\rm Conf}({\mathcal{C}})$, there exist smooth Einstein metrics $(M,g)$ such that $\Phi_{*}(g)=(g_{S},K,[g_{{\mathcal{C}}}])=(\gamma_{S},\kappa,[\gamma_{{\mathcal{C}}}])$.
+However the map $\Phi_{*}$ is of course no longer injective.
+There is a full scalar degree of freedom in the domain $\mathcal{E}$} mapping to the same point in target space.
+Theorem 1.1 shows that the fiber $(\Phi_{*})^{-1}(\gamma_{S},\kappa,[\gamma_{{\mathcal{C}}}])$ in $\mathcal{E}$} is naturally and smoothly identified with $\{\mu\}=C^{\infty}({\mathcal{C}})$.
+This is less clear when $\mu$ is replaced by the mean curvature $H$ of the boundary.
 Note that the boundary data $\mathcal{B}$} is invariant under a restricted group of unimodular or volume-preserving diffeomorphisms. Thus, given $g\in\mathbb{E}^{H}$, let
@@
 We may then form the bundle
-$\widetilde{\mathbb{E}}=\cup_{g\in\mathbb{E}^{H}}{\mathcal{O}}_{\mathrm{Diff}_{0}^{dv_{g}}(M)}.$
+$\widetilde{\mathbb{E}}=\bigcup_{g\in\mathbb{E}^{H}}{\mathcal{O}}_{\mathrm{Diff}_{0}^{dv_{g}}(M)}.$
 If $\widetilde{\pi}:\widetilde{\mathbb{E}}\rightarrow\mathbb{E}^{H}$ is the projection map, then one does have
@@
 $\varphi_{1}^{*}dv_{g_{1}}(dv_{(g_{1})_{{\mathcal{C}}}})^{-2/n}=\varphi_{2}^{*}dv_{g_{2}}(dv_{(g_{2})_{{\mathcal{C}}}})^{-2/n}.$
-We point out that the uniqueness for the gauged map $\Phi^{H}$ in (1.13) cannot be improved to the ungauged map $\Phi$ in (1.5).
+We point out that this uniqueness for the gauged map $\Phi^{H}$ in (1.14) cannot be improved to the ungauged map $\Phi$ in (1.5).
 Thus it is not true that for general $g_{1},g_{2}\in\mathbb{E}$ (not necessarily in harmonic gauge),
@@
 $\Phi$ in (1.5) cannot be one-to-one modulo isometry in $\mathrm{Diff}_{0}(M)$.
-To see this, suppose $g^{1},g^{2}$ are a pair of vacuum Einstein metrics with the same initial data and the same Dirichlet boundary data so that $[g^{1}_{{\mathcal{C}}}]=[g^{2}_{{\mathcal{C}}}]$ and $dv_{g^{1}_{{\mathcal{C}}}}=dv_{g^{2}_{{\mathcal{C}}}}$.
+To see this, suppose $g^{1},g^{2}$ are a pair of vacuum Einstein metrics with the same initial data and the same Dirichlet boundary data so that $g^{1}_{{\mathcal{C}}}]=[g^{2}_{{\mathcal{C}}}]$ and $dv_{g^{1}_{{\mathcal{C}}}}=dv_{g^{2}_{{\mathcal{C}}}}$.
 As noted above, $\mathrm{Diff}_{0}(M)$ acts transitively on the space of bulk volume forms $dv_{g}$ restricted to $\mathcal{C}$}.
@@
 However, this was proved to be false in general, (i.e.
-without further conditions), in [3], [4]; it follows that (1.14) is also false in general.
+without further conditions), in [3], [4], [5]; it follows that (1.15) is also false in general.
 The proof of Theorem 1.1 follows the general lines of the main results in [2, 3].
@@
 Define the Sobolev $H^{s}$ norm on functions on $D$ by
-[…]
+$v||_{H^{s}(D)}^{2}=\sum_{k=0}^{s}\int_{D}|\partial_{D}^{k}v|^{2}dv_{D},$
 where $\partial_{D}^{k}$ consists of all partial coordinate derivatives tangent to $D$ of order $\leq s$.
@@
 Define the stronger $\bar{H}^{s}$ norm by including all space-time derivatives up to order $s$, so that
-[…]
+$v||_{\bar{H}^{s}(D)}^{2}=\sum_{k=0}^{s}\int_{D}|\partial_{M}^{k}v|^{2}dv_{D},$
 where $\partial_{M}$ denotes partial derivatives along all coordinates of $M$ at $D$.
 Finally, for the Cauchy slices $S_{t}$, define the boundary stable $H^{s}$ norm on $S_{t}$ by
-$||v||_{\bar{\mathcal{H}}^{s}(S_{t})}^{2}=||v||_{\bar{H}^{s}(S_{t})}^{2}+||v||_{\bar{H}^{s}({\mathcal{C}}_{t})}^{2}.$
+[…]
 We also recall the Sobolev trace theorem for the restriction map $f:D\rightarrow\mathbb{R}$ to $f|_{\partial D}:\partial D\rightarrow\mathbb{R}$:
-$||f||_{H^{s-1/2}(\partial D)}\leq C||f||_{H^{s}(D)}.$
+[…]
 It will always be assumed that $s$ is sufficiently large, depending on $n$, so that by the Sobolev embedding theorem, $H^{s}\subset C^{2}$.
@@
 Thus (for smooth metrics $g$) the coefficients are smoothly close to constant functions in the rescaled chart,
-(2.5) $||\widetilde{g}-g_{\alpha_{0}}||_{C^{\infty}(U)}\leq\varepsilon=\varepsilon(\lambda,g),$
+(2.5) $\widetilde{g}-g_{\alpha_{0}}||_{C^{\infty}(U)}\leq\varepsilon=\varepsilon(\lambda,g),$
 where $g_{\alpha_{0}}$ is a flat (constant coefficent) Minkowski metric. This is the frozen coefficient approximation.
@@
 Note also that the leading order Sobolev norms scale as follows: for $D\subset S$ or $D\subset{\mathcal{C}}$ of dimension $n$,
-$||f||_{H_{\widetilde{g}}^{s}(D)}^{2}=\lambda^{2s-n}||f||_{H_{g}^{s}(D)}^{2},$
-$||f||_{{\mathcal{N}}_{\widetilde{g}}^{s}(M)}^{2}=\lambda^{2s-n}||f||_{{\mathcal{N}}_{g}^{s}(M)}^{2}.$
+[…]
+[…]
 Also as in [2], $U$ is always assumed to be smoothly embedded in a larger region $\widetilde{U}$,
@@
 $C:{\mathcal{U}}_{c}\times\mathbb{V}_{S}\rightarrow\mathbb{V}_{S},$
-(2.15) $\begin{split}&C(Q,\gamma_{S},\kappa,V_{S},\partial_{t}V_{S},\nu_{S},\gamma_{{\mathcal{C}}},V_{{\mathcal{C}}})=\begin{cases}|\kappa|^{2}-(\mathrm{tr}_{\gamma_{S}}\kappa)^{2}-R_{\gamma_{S}}-\mathrm{tr}_{g_{S}}(Q-\delta^{*}V)+3(Q(\nu_{S},\nu_{S})-\delta^{*}V(\nu_{S},\nu_{S})),\\ {\rm div_{\gamma_{S}}}\kappa-d_{S}(\mathrm{tr}_{\gamma_{S}}\kappa)-Q(\nu_{S},\cdot)+\delta^{*}V(\nu_{S},\cdot).\\ \end{cases}\end{split}$
+[…]
 We then define
@@
 In later sections, we will repeatedly need to solve linear wave equations of the form (2.21), i.e.
-(2.22) $-\tfrac{1}{2}\Box_{g}v+p(v)=\varphi,$
+(2.22) $\tfrac{1}{2}\Box_{g}v+p(v)=\varphi,$
 for given $\varphi$.
@@
 Thus, the strong or boundary stable energy estimate with Dirichlet boundary data states that any (smooth) solution of (2.22) satisfies the bound
-(2.23) $||v||_{\bar{\mathcal{H}}^{s}(S_{t})}^{2}\leq C[||v||_{H^{s}(S_{0})}^{2}+||\partial_{t}v||_{H^{s-1}(S_{0})}^{2}+||v||_{H^{s}({\mathcal{C}}_{t})}^{2}+||\varphi||_{H^{s-1}(M_{t})}^{2}],$
+[…]
 where $t\in[0,1]$ and $C$ is a constant depending only on $g$ and the coefficients of $p$.
@@
 The optimal result in general is:
-(2.24) $|v||_{\bar{\mathcal{H}}^{s}(S_{t})}^{2}\leq C[||v||_{H^{s}(S_{0})}^{2}+||\partial_{t}v||_{H^{s-1}(S_{0})}^{2}+||\nu_{{\mathcal{C}}}(v)||_{H^{s-1/3}({\mathcal{C}}_{t})}^{2}+||\varphi||_{H^{s-1}(M_{t})}^{2}],$
-so there is a ’loss’ of $2/3$ of a derivative. We refer to [9, Ch.8], and in particular to [12, Thm.9] for the proof and further details.
+[…]
+so there is a ’loss’ of $2/3$ of a derivative. We refer to [11, Ch.8], and in particular to [14, Thm.9] for the proof and further details.
 We next discuss the gauge term $V$ and its linearization $V^{\prime}_{h}$. Consider the bulk equation
@@
 The gauge field $V$ satisfies the equation
-(2.26) $-\tfrac{1}{2}(\Box V+\mathrm{Ric}_{g}(V))=\beta(Q).$
+(2.26) $\tfrac{1}{2}(\Box V+\mathrm{Ric}_{g}(V))=\beta(Q).$
 on $M$, where $\Box$ denotes the vector wave operator $\Box=-D^{*}D$ acting on vector fields.
@@
 Similarly, for the linearization $L(h)=F$ as in (2.20), one has
-(2.27) $-\tfrac{1}{2}(\Box_{g}V^{\prime}_{h}+\mathrm{Ric}_{g}(V^{\prime}_{h}))+\tfrac{1}{2}\nabla_{V}V^{\prime}_{h}=\beta(F)+\beta^{\prime}_{h}\mathrm{Ric}+O_{2,1}(V,h),$
+(2.27) $\tfrac{1}{2}(\Box_{g}V^{\prime}_{h}+\mathrm{Ric}_{g}(V^{\prime}_{h}))+\tfrac{1}{2}\nabla_{V}V^{\prime}_{h}=\beta(F)+\beta^{\prime}_{h}\mathrm{Ric}+O_{2,1}(V,h),$
 where $O_{2,1}(V,h)$ is second order in $V$, first order in $h$ and vanishes if $V=0$.
@@
 where $V^{\prime}_{F}=V^{\prime}_{F}({\mathcal{T}})$ is the vector field satisfying
-(2.29) $-\tfrac{1}{2}[\Box V^{\prime}_{F}+\mathrm{Ric}(V^{\prime}_{F})]+\tfrac{1}{2}\nabla_{V}V^{\prime}_{F}=\beta(F),$
+(2.29) $\tfrac{1}{2}[\Box V^{\prime}_{F}+\mathrm{Ric}(V^{\prime}_{F})]+\tfrac{1}{2}\nabla_{V}V^{\prime}_{F}=\beta(F),$
 with $(V^{\prime}_{F})|_{S}=(V^{\prime}_{h})|_{S}$, $(\partial_{t}V^{\prime}_{F})|_{S}=(\partial_{t}V^{\prime}_{h})|_{S}$ along $S$ and $(V^{\prime}_{F})|_{{\mathcal{C}}}=V^{\prime}_{h}|_{{\mathcal{C}}}$ along $\mathcal{C}$}.
@@
 The remaining component $W^{\prime}_{h}$ then satisfies
-(2.30) $-\tfrac{1}{2}[\Box W^{\prime}_{h}+\mathrm{Ric}(W^{\prime}_{h})]+\tfrac{1}{2}\nabla_{V}W^{\prime}_{h}=\beta^{\prime}_{h}\mathrm{Ric}+O_{2,1}(V,h),$
+(2.30) $\tfrac{1}{2}[\Box W^{\prime}_{h}+\mathrm{Ric}(W^{\prime}_{h})]+\tfrac{1}{2}\nabla_{V}W^{\prime}_{h}=\beta^{\prime}_{h}\mathrm{Ric}+O_{2,1}(V,h),$
 with $W^{\prime}_{h}|_{S}=\partial_{t}W^{\prime}_{h}|_{S}=W^{\prime}_{h}|_{{\mathcal{C}}}=0$.
@@
 2.4. Geometry at the corner
-As noted in [2], the conformal factor $\varphi$ of the given boundary data $[g_{{\mathcal{C}}}]$ is determined at the corner $\Sigma$ by the initial data $g_{S}$.
+As noted in [2], the conformal factor $\varphi$ of the given boundary data $g_{{\mathcal{C}}}]$ is determined at the corner $\Sigma$ by the initial data $g_{S}$.
 Consequently, the full metric $g_{\mu\nu}$ at $\Sigma$ is determined at the corner except possibly for the corner angle
@@
 Remark 2.7.
-Let $\sigma$ denote a fixed background metric in the conformal class $[g_{{\mathcal{C}}}]$, so that $g_{{\mathcal{C}}}=\varphi^{2}\sigma$, where $[\sigma]$ is determined by the target data but $\varphi$ is not.
+Let $\sigma$ denote a fixed background metric in the conformal class $g_{{\mathcal{C}}}]$, so that $g_{{\mathcal{C}}}=\varphi^{2}\sigma$, where $\sigma]$ is determined by the target data but $\varphi$ is not.
 As noted above, $\varphi|_{\Sigma}$ is determined by the initial and boundary data.
@@
 The boundary data in the target space ${\mathcal{T}}^{H}$ consists of the data $([g_{{\mathcal{C}}}],\mu_{g},V_{{\mathcal{C}}})$. The linearizations of these terms are given by
-(3.1) $[g_{{\mathcal{C}}}]^{\prime}_{h}=\ring{h}=h^{\mathsf{T}}-\frac{\mathrm{tr}_{{\mathcal{C}}}h^{\mathsf{T}}}{n}g_{{\mathcal{C}}},$
+(3.1) $g_{{\mathcal{C}}}]^{\prime}_{h}=\ring{h}=h^{\mathsf{T}}-\frac{\mathrm{tr}_{{\mathcal{C}}}h^{\mathsf{T}}}{n}g_{{\mathcal{C}}},$
 (3.2) $(\mu_{g})^{\prime}_{h}=\tfrac{1}{2}(h_{\nu\nu}+\tfrac{n-2}{n}\mathrm{tr}_{{\mathcal{C}}}h^{\mathsf{T}})\mu_{g},$
@@
 The function $u$ satisfies the wave equation
-(3.6) $-\tfrac{n-1}{n}\Box_{\mathcal{C}}u-\tfrac{1}{n}R_{\mathcal{C}}u=Y({\mathcal{T}})+E(h),$
+(3.6) $\tfrac{n-1}{n}\Box_{\mathcal{C}}u-\tfrac{1}{n}R_{\mathcal{C}}u=Y({\mathcal{T}})+E(h),$
 along the boundary $\mathcal{C}$}, where
@@
 Theorem 3.2.
-There exists $\varepsilon_{0}>0$ such that, for any smooth metric $g$ on $\widetilde{U}$ which is $C^{\infty}$ $\varepsilon$-close, $\varepsilon\leq\varepsilon_{0}$, to a standard Minkowski corner metric $g_{\alpha_{0}}$ as in (2.5), the linearization $D\Phi_{g}^{H}$ at $g$ satisfies: given any $C^{\infty}$ target data $\tau^{\prime}\in T_{\tau}({\mathcal{T}}^{H})$ on $\widetilde{U}$, there exists a variation $h\in T_{g}Met(\widetilde{U})$, such that
+[…]
 (3.9) $D\Phi_{g}^{H}(h)=\tau^{\prime}.$
@@
 Given this Dirichlet boundary data, solve next the bulk tangential equation
-$-\tfrac{1}{2}\Box_{g}h^{\mathsf{T}}=F^{\mathsf{T}},$
+$\tfrac{1}{2}\Box_{g}h^{\mathsf{T}}=F^{\mathsf{T}},$
 with the given initial conditions on $\widetilde{U}$, as discussed in §2.4 and Lemma 2.1.
 Turning next to $h_{\nu\nu}$, the target data $\mu^{\prime}=\frac{1}{2}(h_{\nu\nu}+\frac{n-2}{n}u)\mu_{g}$ and $u$ above determine the Dirichlet boundary value of $h_{\nu\nu}$ on $\mathcal{C}$} and so we solve the bulk equation
-$-\tfrac{1}{2}\Box_{g}h_{\nu\nu}=F_{\nu\nu},$
+$\tfrac{1}{2}\Box_{g}h_{\nu\nu}=F_{\nu\nu},$
 with these given Dirichlet boundary conditions for $h_{\nu\nu}$ and with the given initial conditions from $\tau^{\prime}$.
 Finally, referring to (3.5), with $\widetilde{Y}$ determined by the target data $\tau^{\prime}$ and setting $\widetilde{E}=0$, the equation (3.5) gives Neumann boundary data $\nabla_{\nu}h(\nu)^{\mathsf{T}}$ for $h(\nu)^{\mathsf{T}}$ and so we solve
-$-\tfrac{1}{2}\Box_{g}h(\nu)^{\mathsf{T}}=F(\nu)^{\mathsf{T}},$
+$\tfrac{1}{2}\Box_{g}h(\nu)^{\mathsf{T}}=F(\nu)^{\mathsf{T}},$
 in $\widetilde{U}$ with this given Neumann boundary data and with initial data given by the target data in $T({\mathcal{T}})$,
 Summing up the solutions above gives a solution $h_{0}$ to
-$-\tfrac{1}{2}\Box_{g}h_{0}=F,$
+$\tfrac{1}{2}\Box_{g}h_{0}=F,$
 in $\widetilde{U}$ with the prescribed initial and boundary conditions from $\tau^{\prime}$.
@@
 Using this, we then solve the bulk equation
-$-\tfrac{1}{2}\Box_{g}E_{1}^{\mathsf{T}}=-P^{\mathsf{T}}(h_{0}),$
+$\tfrac{1}{2}\Box_{g}E_{1}^{\mathsf{T}}=-P^{\mathsf{T}}(h_{0}),$
 in $\widetilde{U}$ with zero initial data and Dirichlet boundary data given by $h^{\mathsf{T}}=\frac{u_{1}}{n}g_{{\mathcal{C}}}$ on $\mathcal{C}$}.
@@
 Next, for $(E_{1})_{\nu\nu}$, setting $\mu_{1}=0$, we use the data $(E_{1})_{\nu\nu}=-\frac{n-2}{n}u_{1}$ as Dirichlet boundary data and solve the IBVP
-$-\tfrac{1}{2}\Box_{g}(E_{1})_{\nu\nu}=-P_{\nu\nu}(h_{0}),$
+$\tfrac{1}{2}\Box_{g}(E_{1})_{\nu\nu}=-P_{\nu\nu}(h_{0}),$
 in $\widetilde{U}$, with zero initial data and with the Dirichlet data given above.
 Finally, solve the IBVP
-$-\tfrac{1}{2}\Box_{g}E_{1}(\nu)^{\mathsf{T}}=-P(\nu)^{\mathsf{T}}(h_{0})$
+$\tfrac{1}{2}\Box_{g}E_{1}(\nu)^{\mathsf{T}}=-P(\nu)^{\mathsf{T}}(h_{0})$
 in $\widetilde{U}$, with zero initial data and with Neumann boundary data given by (3.5) with $\widetilde{Y}=0$ and $\widetilde{E}=\widetilde{E}(h_{0})$.
 Summing these components together then gives $E_{1}$ solving
-$-\tfrac{1}{2}\Box_{g}E_{1}=-P(h_{0}),$
+$\tfrac{1}{2}\Box_{g}E_{1}=-P(h_{0}),$
 in $\widetilde{U}$. The term $E_{1}$ has zero target initial data and zero target boundary data. As noted above, the boundary data for $E_{1}$ is $O(\varepsilon)$ in $C^{\infty}$, so that
@@
 2. Next, let $h_{2}=h_{1}+E_{2}$, where $E_{2}$ is constructed in the same way as $E_{1}$ above, with $E_{1}$ in place of $h_{0}$ and $E_{2}$ in place of $E_{1}$. Thus, $E_{2}$ solves
-$-\tfrac{1}{2}\Box_{g}E_{2}=-P(E_{1}),$
+$\tfrac{1}{2}\Box_{g}E_{2}=-P(E_{1}),$
 in $\widetilde{U}$, with zero target initial data and zero target boundary data.
@@
 in $C^{\infty}$ and hence the sequence $h_{k}=h_{0}+\sum_{1}^{k}E_{i}$ converges to a limit $\bar{h}\in C^{\infty}$. By construction
-$-\tfrac{1}{2}\Box_{g}(h_{0}+\sum E_{i})=F-P(h_{0}+\sum E_{i}),$
+$\tfrac{1}{2}\Box_{g}(h_{0}+\sum E_{i})=F-P(h_{0}+\sum E_{i}),$
 so that
@@
 To address this, from the iterative construction of the component $u$ of $\bar{h}$, it follows by taking limit of equation (3.6) that
-(3.16) $\begin{split}-\tfrac{n-1}{n}\Box_{\mathcal{C}}\bar{u}-\tfrac{1}{n}R_{\mathcal{C}}\bar{u}=&R^{\prime}_{\ring{h}}-\mathrm{tr}(F-\delta^{*}\bar{V}^{\prime})+2(F-\delta^{*}\bar{V}^{\prime})(\nu,\nu)\\ &+2\langle A^{\prime}_{\bar{h}},A\rangle-2H_{\mathcal{C}}H^{\prime}_{\bar{h}}-2\langle\widehat{A}\circ A,\bar{h}\rangle+4\mathrm{Ric}(\nu,\nu^{\prime}_{\bar{h}})+\langle\mathrm{Ric},\bar{h}\rangle\\ &+\mathrm{tr}[(\delta^{*})^{\prime}_{\bar{h}}V]-2[(\delta^{*})^{\prime}_{\bar{h}}V](\nu,\nu),\\ \end{split}$
+[…]
 along $\mathcal{C}$}.
@@
 Define a Sobolev $H^{s}$ norm on the target space data in $T_{\tau}({\mathcal{T}}^{H})$ as follows: for $\tau^{\prime}$ as in (3.8),
-(4.1) $\begin{split}||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}=[||F||_{H^{s}(\widetilde{U})}&+[||\ring{\gamma^{\prime}}||_{H^{s+1}({\mathcal{C}})}+||\mu^{\prime}||_{H^{s+1}({\mathcal{C}})}+||V^{\prime}_{{\mathcal{C}}}||_{H^{s}({\mathcal{C}})}]\\ &+[||\gamma^{\prime}_{S}||_{H^{s+1}(S_{0})}+||\kappa^{\prime}||_{H^{s}(S_{0})}+||\nu^{\prime}||_{H^{s}(S_{0})}+||V^{\prime}_{S}||_{H^{s}(S_{0})}].\end{split}$
+[…]
 On the right in (4.1), the top line consists of bulk and boundary terms, while the bottom line consists of initial data terms.
@@
 For the linearization $D\Phi_{g}^{H}$ of $\Phi^{H}$ in (2.10) with $C^{\infty}$ data localized on $\widetilde{U}\supset U$ as in §2.2, one has the tame energy estimate
-(4.2) $||h||_{{\mathcal{N}}^{s}(S_{t})}\leq C||D\Phi_{g}(h)||_{H^{s}({\mathcal{T}}^{H})},$
+[…]
 where $C$ is a constant depending only on $g$. In particular,
@@
 First decompose $h$ into its components as in (3.14). The bulk equation (2.21) then becomes the following collection of equations:
-(4.3) $-\tfrac{1}{2}\Box_{g}h^{\mathsf{T}}+P^{\mathsf{T}}(h)=F^{\mathsf{T}},$
-(4.4) $-\tfrac{1}{2}\Box_{g}h_{\nu\nu}+P_{\nu\nu}(h)=F_{\nu\nu},$
-(4.5) $-\tfrac{1}{2}\Box_{g}h(\nu)^{\mathsf{T}}+P(\nu)^{\mathsf{T}}(h)=F(\nu)^{\mathsf{T}},$
+(4.3) $\tfrac{1}{2}\Box_{g}h^{\mathsf{T}}+P^{\mathsf{T}}(h)=F^{\mathsf{T}},$
+(4.4) $\tfrac{1}{2}\Box_{g}h_{\nu\nu}+P_{\nu\nu}(h)=F_{\nu\nu},$
+(4.5) $\tfrac{1}{2}\Box_{g}h(\nu)^{\mathsf{T}}+P(\nu)^{\mathsf{T}}(h)=F(\nu)^{\mathsf{T}},$
 By Lemma 2.1, solutions of these equations satisfy the boundary stable energy estimate (2.23) for Dirichlet boundary data.
@@
 Next the evolution or wave equation (3.6) for $u$ along $\mathcal{C}$} gives the following energy estimate for $u$ along $\mathcal{C}$}:
-(4.6) $\begin{split}||u||_{\widetilde{H}^{s}(\Sigma_{t})}\leq C&[||F||_{H^{s-1}({\mathcal{C}}_{t})}+||\ring{h}||_{H^{s+1}({{\mathcal{C}}_{t})}}+||V^{\prime}_{h}||_{\bar{H}^{s}({\mathcal{C}}_{t})}+\varepsilon||h||_{\bar{H}^{s}({\mathcal{C}}_{t})}]\\ +&\ C[||u||_{H^{s}(\Sigma_{0})}+||\partial_{t}u||_{H^{s-1}(\Sigma_{0})}].\end{split}$
+[…]
 Here $\widetilde{H}^{s}(\Sigma_{t})$ denotes the Sobolev norm with both $t$-derivatives as well as derivatives tangent to $\Sigma_{t}$, but no derivatives normal to $\mathcal{C}$}.
@@
 The last line in (4.6) consists of initial data and is bounded by
-$||u||_{H^{s}(\Sigma_{0})}+||\partial_{t}u||_{H^{s-1}(\Sigma_{0})}\leq C[||\gamma^{\prime}_{S}||_{H^{s+1/2}(S_{0})}+||\kappa^{\prime}||_{H^{s-1/2}(S_{0})}],$
+[…]
 where we have used the Sobolev trace theorem and Remark 2.4 and Remark 2.7. Similarly,
-$||F||_{H^{s-1}({\mathcal{C}})}\leq C||F||_{H^{s-1/2}(\widetilde{U})}.$
+[…]
 Next from the gauge wave equation (2.27) and the boundary stable energy estimate with Dirichlet boundary data (2.23) we obtain
-(4.7) $\begin{split}||V^{\prime}_{h}||_{\bar{H}^{s}({\mathcal{C}}_{t})}\leq C&[||F||_{H^{s}(\widetilde{U})}+||V^{\prime}_{{\mathcal{C}}}||_{H^{s}({\mathcal{C}}_{t})}+\varepsilon||h||_{H^{s}(\widetilde{U})}]\\ +&C[||V^{\prime}_{h}||_{H^{s}(S_{0})}+||\partial_{t}V^{\prime}_{h}||_{H^{s-1}(S_{0})}].\end{split}$
+[…]
 Note that this estimate requires an extra derivative on the bulk $F$ term, due to the $\beta(F)$ term in (2.27).
@@
 To obtain a bound on this term, we use the linearized constraint equations (2.13)-(2.14); this gives
-$||\partial_{t}V^{\prime}_{h}||_{H^{s-1}(S_{0})}\leq C[||h_{S}||_{H^{s+1}(S_{0})}+||K^{\prime}_{h}||_{H^{s}(S_{0})}+||F||_{H^{s-1}(S_{0})}+||V^{\prime}_{h}||_{H^{s}(S_{0})}]\leq C||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}.$
+[…]
 Given this control on the Dirichlet boundary value of $h^{\mathsf{T}}$, (4.3) and (2.23) then gives the estimate
-(4.8) $\begin{split}||h^{\mathsf{T}}||_{\bar{\mathcal{H}}^{s}(S_{t})}&\leq C[||h^{\mathsf{T}}||_{H^{s}({\mathcal{C}}_{t})}+\varepsilon||h||_{H^{s}(\widetilde{U}_{t})}+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}]\\ &\leq C[\int_{0}^{t}[||\ring{h}||_{H^{s+1}({{\mathcal{C}}_{t^{\prime}})}}+||V^{\prime}_{{\mathcal{C}}}||_{H^{s}({\mathcal{C}}_{t^{\prime}})}.+\varepsilon||h||_{\bar{H}^{s}({\mathcal{C}}_{t^{\prime}})}]dt^{\prime}+\varepsilon||h||_{H^{s}(\widetilde{U}_{t})}+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}],\\ \end{split}$
+[…]
 where the last inequality comes from integration of (4.6)-(4.7) over $t^{\prime}\in[0,t]$. Taking the maximum over $t\in I$ then gives
-(4.9) $\max_{t\in[0,1]}||h^{\mathsf{T}}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C[||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}+\varepsilon||h||_{\widehat{H}^{s}(\widetilde{U})}],$
-where $||h||_{\widehat{H}^{s}(\widetilde{U})}:=||h||_{H^{s}(\widetilde{U})}+||h||_{\bar{H}^{s}({\mathcal{C}})}$.
+[…]
+[…]
 Similarly, the Dirichlet energy estimate for $h_{\nu\nu}$ gives
-$\max_{t\in[0,1]}||h_{\nu\nu}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C[||h_{\nu\nu}||_{H^{s}({{\mathcal{C}})}}+\varepsilon||h||_{H^{s}(\widetilde{U})}+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}],$
+[…]
 and so using the definition of $\mu^{\prime}_{h}$, $h_{\nu\nu}=\mu^{\prime}_{h}-\frac{n-2}{n}u$, together with (4.9) gives
-(4.10) $\max_{t\in[0,1]}||h_{\nu\nu}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C[||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}+\varepsilon||h||_{\widehat{H}^{s}(\widetilde{U})}],$
+[…]
 Finally for $h(\nu)^{\mathsf{T}}$, from the Neumann boundary estimate (2.24), we have
-$||h(\nu)^{\mathsf{T}}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C[||\nabla_{\nu}h(\nu)^{\mathsf{T}}||_{H^{s-1/3}({{\mathcal{C}}_{t})}}+\varepsilon||h||_{H^{s}(\widetilde{U}_{t})}+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}].$
+[…]
 Examination of the form of the Neumann boundary data in (3.3) gives
-(4.11) $\begin{split}||h(\nu)^{\mathsf{T}}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq&C[||\ring{h}||_{H^{s+2/3}({\mathcal{C}}_{t})}+||\mu^{\prime}_{h}||_{H^{s+2/3}({\mathcal{C}}_{t})}\\ &+||V^{\prime}_{{\mathcal{C}}}||_{H^{s-1/3}({\mathcal{C}})}+\varepsilon(||h||_{H^{s-1/3}({\mathcal{C}}_{t})}+||h||_{H^{s}(\widetilde{U}_{t})})+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}].\end{split}$
-Summing up the estimates (4.6)-(4.11), by choosing $\varepsilon$ sufficiently small depending on $g$, one may absorb the $\varepsilon||h||_{\bar{H}^{s}(\widetilde{U})}$ terms on the right side of these equations into the total sum on the left, using the fact that $|h||_{\bar{H}^{s}(\widetilde{U})}\leq\max_{t\in[0,1]}|h||_{\bar{\mathcal{H}}^{s}(S_{t})}$, giving then
-(4.12) $\max_{t\in[0,1]}||h||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}.$
+[…]
+[…]
+[…]
 We note that the estimate (4.11) is the key ingredient in establishing the bound (4.12) in comparison with the work in [2, 3].
@@
 so that (4.12) gives in fact
-(4.14) $||h||_{{\mathcal{N}}^{s}}^{2}\leq C||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}.$
+[…]
 This completes the proof of Theorem 4.1.
@@
 Moreover, any $h$ satisfying (5.1) satisfies the tame energy estimate
-(5.2) $||h||_{{\mathcal{N}}^{s}}\leq C||D\Phi_{g}^{H}(h)||_{H^{s}(T^{H})},$
+[…]
 where $C$ depends only on $g$.
@@
 $L(h_{0})=\rho_{0}F\ \mbox{ in }\widetilde{U}_{0},\ \ (g_{S},K,\nu_{S},V_{g})^{\prime}_{h}=\rho_{0}(\gamma^{\prime},\kappa^{\prime},\nu^{\prime},V_{S}^{\prime})\ \mbox{ on }S_{0}.$
-Then again $h_{0}$ has compact support in $\widetilde{U}_{0}$ for some time interval $[0,t_{0}]$ and as above extends to $M_{t_{0}}$.
+Then again $h_{0}$ has compact support in $\widetilde{U}_{0}$ for some time interval $0,t_{0}]$ and as above extends to $M_{t_{0}}$.
 We relabel so that $t_{0}$ is a common time interval for all $h_{i}$, $i\geq 0$.
@@
 However, by (2.5), the metrics $\widetilde{g}_{i}$ satisfy
-$||\widetilde{g}_{2}-\widetilde{g}_{1}||_{C^{\infty}}\leq\varepsilon,$
+$\widetilde{g}_{2}-\widetilde{g}_{1}||_{C^{\infty}}\leq\varepsilon,$
 in $U$, for a prescribed $\varepsilon>0$. It then follows from the local uniqueness in the Nash-Moser theorem above that
@@
 The group $\mathrm{Diff}_{0}(M)/\mathrm{Diff}_{1}(M)$ is a smooth tame Frechet Lie group (cf.
-[7, I.4.6]) which acts smoothly and freely on $\mathbb{E}^{H}$.
+[8, I.4.6]) which acts smoothly and freely on $\mathbb{E}^{H}$.
 It then follows from standard theory, cf.
-[7, Cor.III.2.5.2] that the quotient $\mathbb{E}^{H}/(\mathrm{Diff}_{0}(M)/\mathrm{Diff}_{1}(M))$ is a smooth, tame Frechet manifold.
+[8, Cor.III.2.5.2] that the quotient $\mathbb{E}^{H}/(\mathrm{Diff}_{0}(M)/\mathrm{Diff}_{1}(M))$ is a smooth, tame Frechet manifold.
 As noted in §1,
@@
 The energy estimates in (5.2) hold on the $H^{s}$ completions of the domain and target spaces for suitable values of $s$, giving the existence of a tame inverse for the derivative $D\Phi^{H}$.
-One may then apply the finite derivative version of the Nash-Moser inverse function theorem due to Zehnder [13], or [14, Thm.6.3.3]; in the notation used there, the loss of derivative is $\gamma=3$.
-However, we will not carry out here the details that the hypotheses of [14, Thm.6.3.3] do hold.
+One may then apply the finite derivative version of the Nash-Moser inverse function theorem due to Zehnder [15], or [16, Thm.6.3.3]; in the notation used there, the loss of derivative is $\gamma=3$.
+However, we will not carry out here the details that the hypotheses of [16, Thm.6.3.3] do hold.
 Remark 6.3.